### v2

# Min-K%++: Improved Baseline for Detecting Pre-Training Data from Large Language Models

Abstract
The problem of pre-training data detection for large language models (LLMs) has received growing attention due to its implications in critical issues like copyright violation and test data contamination. A common intuition for this problem is to identify training data by checking if the input comes from a mode of the LLM’s distribution. However, existing approaches, including the state-of-the-art Min-K%, often use zeroth-order signals for detection, which are less robust in determining local maxima than second-order statistics. In this work, we propose a novel methodology Min-K%++ for pre-training data detection that measures how sharply peaked the likelihood is around the input, a measurement analogous to the curvature of continuous distribution. Our method is theoretically motivated by the observation that maximum likelihood training implicitly optimizes the trace of the Hessian matrix of likelihood through score matching. Empirically, the proposed method achieves new SOTA performance across multiple settings. On the WikiMIA benchmark, Min-K%++ outperforms the runner-up by 6.2% to 10.5% in detection AUROC averaged over five models. On the more challenging MIMIR benchmark, it consistently improves upon reference-free methods while performing on par with reference-based method that requires an extra reference model.
1 Introduction
Data is one of the most important factors for the success of large language models (LLMs). As the training corpus grows in scale, it has increasing tendency to be held in-house as proprietary data instead of being publicly disclosed [32, 1]. However, for large-scale training corpora that consist of up to trillions of tokens [9], the sheer volume of the training corpus can lead to unintended negative consequences. For example, memorized private information is vulnerable to data extraction [4], and memorized copyrighted contents (e.g., books and news articles) may violate the rights of content creators [14, 18]. Furthermore, it becomes increasingly likely that evaluation data is exposed at training time, bringing the faithfulness and effectiveness of evaluation benchmarks into question [26].
For these reasons, there has been growing interest in effective pre-training data detection strategies. Pre-training data detection can be considered a special case of Membership Inference Attack (MIA) [29]: the goal is to infer whether a given input has been used for training a target LLM (see Figure 1 left for illustration). Due to characteristics of pre-training corpora and training characteristics of LLMs [28, 10], this problem has been shown to be much more challenging than conventional MIA settings (see Section 5 for details).
In this work, we propose a novel methodology, Min-K%++, for pre-training data detection. The motivation lies in that existing methods (e.g., the state-of-the-art, Min-K[28]) often measure zeroth-order signals which could be less robust than second-order statistics in detecting training data (Section 3.1). Then, by connecting maximum likelihood training with score matching [17], we uncover that maximum likelihood training implicitly minimizes the trace of the Hessian matrix of likelihood over the training data (Section 3.2). This theoretical observation pinpoints a concrete second-order indicator for training data detection under a continuous distribution. Finally, we translate our insights into a practical and effective scoring mechanism that works for the discrete (categorical) distribution modeled by LLMs (Section 3.3). The core idea is to examine how sharply the likelihood is around the input, leading to an analogous measurement to the curvature of continuous distribution.
Figure 1: Left: We study the pre-training data detection problem for LLMs [28, 10]. Given the grey-box access to the target LLM (only the logits, token probabilities, and loss are available), the goal is to design a scoring function that yields a score which (upon thresholding) best separates training and non-training text. Right: ROC (receiver operating characteristic) curves of various methods on WikiMIA benchmark [28]. The AUROC (area under the ROC curve) score is denoted in the legend. Our Min-K%++ improves upon existing approaches by large margin.
Empirically, we perform extensive experiments to verify the proposed method on two established benchmarks, WikiMIA [28] and MIMIR [10]. Our experiments consider up to 5 families of models, including LLaMA [31], Pythia [2], and the new architecture Mamba [15]. Figure 1 right showcases the performance comparison between our method and previous ones. Notably, Min-K%++ on average leads to 6.2% to 10.5% absolute increases over the runner-up Min-Kon WikiMIA. On the more challenging MIMIR benchmark which consists of 7 subsets from various domains of the Pile dataset [13], our method still consistently outperforms Min-Kand is on par with the reference-based method [4]. Importantly, unlike reference-based methods which require another LLM to calibrate the likelihood, Min-K%++ is applied to a standalone target LLM. We also introduce and study an online detection setting that simulates “detect-while-generating”; the proposed Min-K%++ again performs the best. Lastly, Min-K%++ can also be interpreted from a calibration perspective: with ablation study we show that both calibration factors in the formulation contribute to the overall high performance.
2 Background
In this section, we first cover the problem statement of pre-training data detection defined by prior works [29, 28, 10]. Then, we go over the formulation of Min-K[28] as a representative method (due to its superior performance) to give a sense of how existing approaches handle the task.
Problem statement. Pre-training data detection is an instance of Membership Inference Attack (MIA) [29]. Formally, given 1) a data instance $x$ and 2) a pre-trained auto-regressive LLM $\mathcal{M}$ that is trained on a dataset $\mathcal{D}$, the goal is to infer whether $x\in\mathcal{D}$ or not (i.e., $x$ is training data or non-training data). The approach to detection leverages a scoring function $s(x;\mathcal{M})$ that computes a score for each input. A threshold is then applied to the score to yield a binary prediction:
$\text{prediction}(x,\mathcal{M})=\begin{cases}1~{}(x\in\mathcal{D}),&s(x;\mathcal{M})\geq\lambda\\ 0~{}(x\notin\mathcal{D}),&s(x;\mathcal{M})<\lambda\end{cases},$ (1)
where $\lambda$ is a case-dependent threshold. Following the established standard [21, 28, 10], we consider grey-box access to the target model $\mathcal{M}$, meaning that one can only access the output statistics including the loss value, logits, and token probabilities. Additional information such as the model weights and gradients are not available. In summary, the key to the pre-training data detection is to design an appropriate scoring function that best separates training data from non-training data.
Min-Kmethod [28]. Recall that an auto-regressive LLM models the (conditional) categorical distribution of the next token given a prefix (i.e., all preceding tokens). Concretely, given a text input $x$ with $T$ tokens $(x_{1},x_{2},...,x_{T})$, we can obtain the probability $p(x_{t}|x_{<t})$ for each token $x_{t}$ given the prefix $x_{<t}$. [^1] For simplicity, from now on we use $(x_{<t},x_{t})$ to represent the token sequence $x_{1},x_{2},...,x_{t}$. Basing itself on a simple intuition that training texts are less likely to include low-probability “outlier” tokens, Min-Kdirectly measures the log probability of each token as the score with Equation 2. To aggregate the score for the full sentence, Equation 3 picks the $k$of tokens that have the lowest log probability and computes the average log probability over the selected tokens, with $k\in(0,100]$ being a hyperparameter. Formally,
[…]
3 Min-K%++
In this section we introduce our methodology, called Min-K%++. We start by analyzing shortcomings of existing methods, where we point out that the zeroth-order signals measured by them may not be robust for detecting training data; instead, we argue that second-order statistic serves as a more reliable indicator. Then, by connecting the maximum likelihood training with score matching [17], we present a theoretical insight showing that the trace of the Hessian matrix of likelihood (which is second-order) is implicitly optimized on the training data. Finally, translating our insights into a practical solution, we formulate the proposed Min-K%++ which works for the discrete distribution of LLMs and measures an analogy of the curvature of continuous distribution.
3.1 Motivation: zeroth-order v.s. second-order
Intuitively, Min-Kand many existing methods try to identify training data by determining if the input comes from one of the modes of the probability distribution characterised by the model [28, 4]. However, they typically evaluate zeroth-order signals which are less robust than second-order statistics in telling whether the data forms a local maximum, a fact that we know from the derivative test [33]. Taking Min-Kas a representative example, it measures the absolute token probability as in Equation 2 which is only good for finding global maximum rather than local maximum. Later we will present a specific example to showcase such insufficiency.
3.2 Theoretical foundation: Hessian trace is optimized during maximum likelihood training
Pursuing along our motivation, we now wonder if there is a concrete second-order statistic that can help identify training data. Specifically, our exploration in this section features the key theoretical insight of this work, where we uncover that the trace of the Hessian matrix of likelihood (w.r.t. the input) is implicitly optimized through the course of maximum likelihood training. This observation lays solid foundation for later designing our methodology.
Our idea here is to ask an insightful question that, to our knowledge, previous works fail to touch on. That is, what effects or “footprints” do training samples leave to the model after training? To answer the question, first recall that (auto-regressive) LLMs are trained by maximizing the likelihood of each token $x_{t}$ conditioned on its prefix $x_{<t}$, i.e., $\log p(x_{t}|x_{<t})$. Meanwhile, for clarity, let us generalize from LLMs to the general maximum likelihood training for now.
Earlier, Koehler et al. [19] have proved that implicit score matching (ISM) [17] is a relaxation (within a multiplicative factor) of maximum likelihood estimation. Consequently, one can interpret maximum likelihood training loss through the lens of score matching and reformulate it with ISM [17] as
$\frac{1}{N}\sum_{\bm{x}}\biggl{[}\frac{1}{2}||\psi(\bm{x})||^{2}+\underbrace{\sum_{i=1}^{d}\frac{\partial\psi_{i}(\bm{x})}{\partial\bm{x}^{i}}}_{\text{Hessian trace}}\biggr{]},$ (4)
where $\psi(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}}$ is the score function [17] for the $d$-dimensional input $\bm{x}$, $\psi_{i}(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}^{i}}$, $\bm{x}^{i}$ is the element at the $i$-th dimension of $\bm{x}$, and $N$ is the number of training samples. Without ambiguity, we are omitting the model parameters in the formulation of $\psi$ for brevity.
Remark. Notice that the second term in the summation in Equation 4 is the trace of the Hessian matrix of log likelihood. Therefore, the key takeaway here is that (for continuous distributions) maximum likelihood training implicitly minimizes the Hessian trace of likelihood. More specifically, training data will tend to have smaller Hessian trace than non-training data, pinpointing a second-order signal that can provide accurate information for detecting training instances. This observation also in general applies to LLMs (as they are trained with maximum likelihood estimation), with a small caveat being that LLMs model a discrete (categorical) distribution instead of a continuous distribution. We address this caveat when formulating our method, which is discussed in the next section.
3.3 Formulating Min-K%++
Now that we have identified that the second-order Hessian trace is optimized over training data and thus encodes rich information for training data detection, we proceed to actually designing our scoring mechanism. Note, naïvely measuring the Hessian trace cannot work here because 1) the derivative is not well-defined for discrete inputs (i.e., tokens) of LLMs, and 2) even if it is somehow defined (e.g., defining the input as token embedding), computing gradients violates the grey-box access constraint.
Instead, our core idea is to measure how sharply peaked the likelihood is around an input, which takes inspiration from that Hessian matrix describes the curvature of a (continuous) function [35]. Concretely, we propose the following formulation for our Min-K%++:
[…]
Here, $\mu_{x_{<t}}=\mathbb{E}_{z\sim p(\cdot|x_{<t})}[\log p(z|x_{<t})]$ is the expectation of the next token’s log probability over the vocabulary of the model given the prefix $x_{<t}$, and $\sigma_{x_{<t}}=\sqrt{\mathbb{E}_{z\sim p(\cdot|x_{<t})}[(\log p(z|x_{<t})-\mu_{x_{<t}})^{2}]}$ is the standard deviation. In practice, both terms can be computed analytically since the categorical distribution $p(\cdot|x_{<t})$ is encoded by the output logits of the model, which we have access to. Upon the token-wise score computed by Equation 5, we follow the same process as Min-Kto aggregate the sentence-wise score in Equation 6, i.e., we select the $k$of the tokens with the minimum score and compute the average score over them.
Interpretation 1: Measuring how sharply peaked the likelihood is. By doing this, our method puts up a measurement for discrete space that is analogous to the curvature of a continuous distribution. Specifically, the numerator of Equation 5 compares the likelihood $\log p(x_{t}|x_{<t})$, which corresponds to the ground-truth next token $x_{t}$, with the expected likelihood $\mu_{x_{<t}}$ computed over all the candidate tokens within the vocabulary. The larger $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$ is, the more probability mass gets assigned to the input token $x_{t}$ compared to other plausible tokens, meaning that $x_{t}$ tend to form a mode given $x_{<t}$ according to the model’s distribution (thus $(x_{<t},x_{t})$ is more likely to be training data).
To see this perspective in a more straightforward way, let us consider a conceptual example depicted in Figure 2, where we show two instances of the distribution $p(\cdot|x_{<t})$ over a vocabulary of size 10. In the two cases (Figure 2 (a) and (b)), both $p(x_{t}|x_{<t})$ is 0.2, meaning that Min-Kwith this zeroth-order information will assign exactly the same score to the token $x_{t}$ and give the same judgement in both cases (e.g., that the sequence $(x_{<t},x_{t})$ was not seen in the training). However, inspecting them with the idea of our Min-K%++ reveals their distinct characteristics:
  • In Figure 2 (a), one can notice that the distribution is actually peaked at $x_{t}$ (the bar with the hatch pattern ‘//’), indicating that the current sequence $(x_{<t},x_{t})$ forms a mode and is very likely to have appeared in the training.
  • In Figure 2 (b), the probability distribution is not peaked at $x_{t}$ as there is another token that gets assigned much larger probability mass. This means that $x_{t}$ is less likely to follow the prefix $x_{<t}$ according to the learned distribution of the model, which should be seen as an indicator that the sequence $(x_{<t},x_{t})$ was not seen in the training.
The above reasoning can be concisely reflected by comparing $\log p(x_{t}|x_{<t})$ and $\mu_{x_{<t}}$, which are marked by the two dashed lines in Figure 2. Accordingly, it can be seen that Figure 2 (a) and (b) exhibit positive and negative $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$, respectively, which successfully separate them apart. In addition to this high-level understanding, in Appendix A we also analytically show that the numerator shares similar form to an approximation of the Hessian trace, which further validates its design.
Interpretation 2: A calibration perspective. Reference-based methods [4] show superior performance as membership inference attacks by calibrating the (sentence-level) likelihood with certain references, such as the likelihood of the same input on a smaller LLM, the Zlib entropy, or the likelihood of lowercased text. In this regard, Min-K%++ can be seen as calibrating the token likelihood with two calibration factors, $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. The unique advantage of our method lies in that instead of using some external references (e.g., an extra LLM), we leverage the statistics that can be readily computed with the target model only. The effect of $\mu_{x_{<t}}$ has been thoroughly discussed. The inclusion of $\sigma_{x_{<t}}$ is inspired by temperature scaling [16], a technique for calibrating the prediction confidence of neural networks. It scales the model output by a constant and has been shown to benefit tasks like Out-of-Distribution detection [20, 39]. Instead of using a constant, here we use a dynamic factor $\sigma_{x_{<t}}$ that is adaptive to each input and find that it leads to extra performance boosts. Later we will demonstrate that both factors contribute to the success of Min-K%++ through ablation study.
Figure 2: A conceptual example showcasing the idea of Min-K%++. (a) and (b) show two types of categorical distribution over the vocabulary. The bars with the hatch pattern (‘//’) correspond to the target token $x_{t}$. Min-K%++ measures how sharply peaked the likelihood is at the ground-truth next token $x_{t}$ by comparing $\log p(x_{t}|x_{<t})$ with $\mu_{x_{<t}}$.
4 Experiments
We conduct extensive experiments to validate the empirical efficacy of the proposed Min-K%++. We first focus on two established benchmarks for pre-training data detection. Following that, ablation studies are presented and discussed. Lastly, we investigate a new online detection setting.
4.1 Setup
Benchmarks. We focus on two benchmarks (and the only two to our knowledge) for pre-training data detection, WikiMIA [28] and MIMIR [10]. WikiMIA is constructed by the authors of Min-Kand is the first benchmark for pre-training data detection. It consists of texts from Wikipedia events, and the training v.s. non-training data is determined by the timestamp (i.e., all events that happen after the release date of a LLM are guaranteed to be non-training data for that model). WikiMIA specifically groups data into splits according to the sentence length, intending to provide a fine-grained evaluation and isolate the effect of input length. It also considers two settings: original and paraphrased. The former assesses the detection of verbatim training texts, while the latter paraphrases the training texts (using ChatGPT) and evaluates on paraphrased inputs.
MIMIR [10] is built upon the Pile dataset [13], where training samples and non-training samples are drawn from the train and test split, respectively. MIMIR is found to be much more challenging than WikiMIA since the training and non-training texts are from the same dataset and thus have minimal distribution shifts and temporal discrepancy [10].
Baselines. We consider 6 representative and state-of-the-art methods as our baselines, which are also featured in the benchmarking work of Duan et al. [10]. Except Min-K%, all methods consider sentence-level probability. Loss method [37] is a general technique that directly takes the loss as the score for detection. Reference method [4] (Ref) uses a reference LLM to calibrate the likelihood of the input. Zlib and Lowercase method [4] are also reference-based, except that they use zlib compression entropy and the likelihood of lowercased text as the reference, respectively. Neighbor method [21] perturbs the input sentence with masked language models to create “neighbors” and calibrate the loss of the input sentence with the average loss of the neighbor sentences. Lastly, as discussed earlier, Min-K[28] examines the token probability and aggregate the score through min-$k$selection; it is currently the best-performing method on WikiMIA.
Determining the $k$%. Our Min-K%++, just like Min-K%, selects the $k$of tokens with the minimum score to compute the overall score for the sentence. As existing benchmarks do not provide validation data, we sweep over 10, 20, …, 100 for $k$ and report the best result for both Min-Kand Min-K%++ for fair comparison. Later we show that Min-K%++ is robust to this hyperparameter in ablation study.
Models. WikiMIA is applicable to a wide range of models since Wikipedia dumps are often included into the training corpus of many LLMs. Specifically, we consider the suite of Pythia models [2] (2.8B, 6.9B, 12B), GPT-NeoX [3] (20B), LLaMA [31] (13B, 30B, 65B), and OPT [40] (66B). We further include a new family of architectures, Mamba [15]. When a reference model is needed, following Shi et al. [28] we use the smaller version correspondingly, e.g., LLaMA-7B for LLaMA models and Pythia-70M for Pythia models. MIMIR is applicable to models that are trained on Pile. To be consistent with Duan et al. [10], we focus on Pythia models (160M, 1.4B, 2.8B, 6.9B, 12B).
Metrics. As a binary classification problem, the detection performance can be evaluated with the AUROC score (area under the receiver operating characteristic curve) [4, 28, 10]. We define training data as “positive” and non-training data as “negative”. AUROC is threshold-independent and can be interpreted as the the probability that a positive instance has higher score than a negative instance according to the detector. Therefore, the higher the better, and the random-guessing baseline is 50%. While we use AUROC as the main metric, we also report True Positive Rate (TPR) at low False Positive Rate (FPR) which measures detection rate at a meaningful threshold.
Compute. All experiments are run with a single Tesla A100 80 GB GPU. When the model is large, we use either half-precision floating point or 8-bit quantization to fit the model into GPU memory.
Table 1: AUROC results on WikiMIA benchmark [28]. Ori. and Para. denote the original and paraphrased settings, respectively. Bolded number shows the best result within each column across all methods. The proposed Min-K%++ leads to remarkable improvements over existing methods in most settings.
Mamba-1.4B Pythia-6.9B LLaMA-13B LLaMA-30B LLaMA-65B Average
Len. Method Ori. Para. Ori. Para. Ori. Para. Ori. Para. Ori. Para. Ori. Para.
32 Loss [37] 61.0 61.4 63.8 64.1 67.5 68.0 69.4 70.2 70.7 71.8 66.5 67.1
Ref [4] 62.2 62.3 63.6 63.5 57.9 56.2 63.5 62.4 68.8 68.2 63.2 62.5
Lowercase [4] 60.9 60.6 62.2 61.7 64.0 63.2 64.1 61.2 66.5 64.8 63.5 62.3
Zlib [4] 61.9 62.3 64.3 64.2 67.8 68.3 69.8 70.4 71.1 72.0 67.0 67.4
Neighbor [21] 64.1 63.6 65.8 65.5 65.8 65.0 67.6 66.3 69.6 68.7 66.6 65.8
Min-K[28] 63.2 62.9 66.3 65.2 68.0 68.4 70.1 70.7 71.3 72.2 67.8 67.9
Min-K%++ 66.8 66.1 70.3 68.0 84.8 82.7 84.3 81.2 85.1 81.4 78.3 75.9
64 Loss [37] 58.2 56.4 60.7 59.3 63.6 63.1 66.2 65.5 67.9 67.7 63.3 62.4
Ref [4] 60.6 59.6 62.4 62.9 63.4 60.9 69.0 65.4 73.4 71.0 65.8 63.9
Lowercase [4] 57.0 57.0 58.2 57.7 62.0 61.0 62.1 59.8 64.5 61.9 60.8 59.5
Zlib [4] 60.4 59.1 62.6 61.6 65.3 65.3 67.5 67.4 69.1 69.3 65.0 64.5
Neighbor [21] 60.6 60.6 63.2 63.1 64.1 64.7 67.1 66.7 69.6 69.5 64.9 64.9
Min-K[28] 62.2 58.0 65.0 61.1 66.0 64.0 68.5 65.7 69.8 67.9 66.3 63.3
Min-K%++ 67.2 63.3 71.6 64.8 85.7 78.8 84.7 74.9 83.8 74.0 78.6 71.2
128 Loss [37] 63.3 62.7 65.1 64.7 67.8 67.2 70.3 69.2 70.7 70.2 67.4 66.8
Ref [4] 62.0 61.1 63.3 62.9 62.6 59.7 71.9 70.0 73.7 72.0 66.7 65.1
Lowercase [4] 58.5 57.7 60.5 60.0 60.6 56.4 59.1 55.4 63.3 60.1 60.4 57.9
Zlib [4] 65.6 65.3 67.6 67.4 69.7 69.6 71.8 71.5 72.1 72.1 69.4 69.2
Neighbor [21] 64.8 62.6 67.5 64.3 68.3 64.0 72.2 67.2 73.7 70.3 69.3 65.7
Min-K[28] 66.8 64.5 69.5 67.0 71.5 68.7 73.9 70.2 73.6 70.8 71.0 68.2
Min-K%++ 68.8 65.6 70.7 66.8 83.9 76.2 82.6 73.8 80.0 70.7 77.2 70.6
Table 2: AUROC results on the challenging MIMIR benchmark [10] with a suite of Pythia models [2]. In each column, the best result across all methods is bolded, with the runner-up underlined. Ref relies on a reference LLM. Neighbor induces significant extra computational cost than others (25×25$\times$25 × in this case), for which reason we don’t run on the 12B model. Despite not requiring a reference model like the Ref method does, our Min-K%++ consistently achieves superior or comparable performance.
Wikipedia Github Pile CC PubMed Central
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [37] 50.2 51.3 51.8 52.8 53.5 65.7 69.8 71.3 73.0 74.0 49.6 50.0 50.1 50.7 51.1 49.9 49.8 49.9 50.6 51.3
Ref [4] 51.2 55.2 58.1 61.8 63.9 63.9 67.1 65.3 64.4 63.0 49.2 52.2 53.7 54.9 56.7 51.3 53.1 53.7 55.9 58.2
Zlib [4] 51.1 52.0 52.4 53.5 54.3 67.4 71.0 72.3 73.9 74.8 49.6 50.1 50.3 50.8 51.1 49.9 50.0 50.1 50.6 51.2
Neighbor [21] 50.7 51.7 52.2 53.2 / 65.3 69.4 70.5 72.1 / 49.6 50.0 50.1 50.8 / 47.9 49.1 49.7 50.1 /
Min-K[28] 50.2 51.3 51.8 53.6 54.4 65.7 69.9 71.4 73.2 74.3 50.3 51.0 50.8 51.5 51.7 50.6 50.3 50.5 51.2 52.3
Min-K%++ 49.7 53.7 55.1 58.0 61.1 64.8 69.6 70.9 72.8 74.2 50.6 51.0 51.0 53.0 53.5 50.6 51.4 52.4 54.2 55.4
ArXiv DM Mathematics HackerNews Average
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [37] 51.0 51.5 51.9 52.9 53.4 48.8 48.5 48.4 48.5 48.5 49.4 50.5 51.3 52.1 52.8 52.1 53.1 53.5 54.4 54.9
Ref [4] 49.4 51.5 53.1 55.8 57.5 51.1 51.1 50.5 51.1 50.9 49.1 52.2 55.1 57.9 60.6 52.2 54.6 55.6 57.4 58.7
Zlib [4] 50.1 50.9 51.3 52.2 52.7 48.1 48.2 48.0 48.1 48.1 49.7 50.3 50.8 51.2 51.7 52.3 53.2 53.6 54.3 54.8
Neighbor [21] 50.7 51.4 51.8 52.2 / 49.0 47.0 46.8 46.6 / 50.9 51.7 51.5 51.9 / 52.0 52.9 53.2 53.8 /
Min-K[28] 51.0 51.7 52.5 53.6 54.6 49.4 49.7 49.5 49.6 49.7 50.9 51.3 52.6 53.6 54.6 52.6 53.6 54.2 55.2 55.9
Min-K%++ 50.1 51.1 53.7 55.2 58.0 50.5 50.9 51.7 51.6 51.9 50.7 51.3 52.6 54.5 56.5 52.4 54.1 55.3 57.0 58.7
4.2 Results
WikiMIA results. Table 1 shows major results in terms of AUROC; for results on more models and TPR numbers, please see Appendix B Tables 5, 6, 7 and 8. We remark that Min-K%++ achieves significant improvements over existing methods. In the original setting, Min-K%++ on average outperforms the runner-up Min-Kby {10.5%, 12.3%, 6.2%} with inputs of length {32, 64, 128}, respectively. In the paraphrased setting, Min-K%++ is also the best-performing approach on average in all cases.
Min-K%++ is consistent across models and input lengths. It can be noticed that Min-K%++’s superior results are agnostic to models: besides transformer-based LLMs, Min-K%++ also decently outperforms others on the new state space-based architecture, Mamba. In terms of input length, Shi et al. [28] identify that short inputs are more challenging than longer inputs. While this is indeed the case for Min-K%, which yields {62.2%, 66.8%} AUROC on {64, 128}-length inputs with Mamba-1.4B (a 4.6% decrease when changing to shorter inputs), Min-K%++ achieves a much more consistent performances of {67.2%, 68.8%} (a mere decrease of 1.6%). Both observations demonstrate the robustness and generality of our proposed method.
MIMIR results. Table 2 shows the AUROC results; see Appendix B Table 9 for TPR results. Most numbers are taken from those reported by Duan et al. [10]. MIMIR is extremely challenging in that the training and non-training texts are maximally similar to each other since they are drawn from the same dataset. Nonetheless, Min-K%++ still improves upon existing methods in most cases. Averaged over 7 subsets, Min-K%++’s relative AUROC w.r.t. the powerful Min-Kis {–0.2%, +0.5%, +1.1%, +1.8%, +2.8%} on Pythia model with {160M, 1.4B, 2.8B, 6.9B, 12B} parameters, respectively. Extrapolating this trend, we anticipate Min-K%++’s effectiveness to be even more obvious with larger models. The most powerful method on MIMIR is the Reference approach [4] (Ref), the results of which are obtained by exhaustively trying out 8 different LLMs as the reference model and picking the best one [10] (note that inferencing with an extra LLM is expensive and may not always be feasible). In contrast, our Min-K%++ does not rely on a reference model, yet provides competitive performance that is on par with Ref. Meanwhile, we remark that Min-K%++ achieves new SOTA results among all the 5 reference-free methods out of the 6 total baselines. Also, in Appendix B Table 9, Min-K%++ outperforms all baselines when it comes to TPR performance.
4.3 Ablation study
We focus on WikiMIA with LLaMA-13B model for ablation study.
Ablation on the hyperparameter $k$%. $k$ determines what percent of tokens with minimum scores are chosen to compute the final score. From Table 3, it is obvious that Min-K%++ is robust to the choice of $k$, with the best and the worst result being 84.8% and 82.1% (a variation of 2.7%), respectively. Min-Kis a little more sensitive to $k$: the variation between the best (68.0%) and the worst result (63.6%) is 4.4%, slightly larger than that of Min-K%++.
Decomposing the contribution of calibration factors. Recall that our method can be interpreted as calibrating the log probability $\log p(x_{t}|x_{<t})$ with two calibrationg factors $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ (Equation 5). In Table 3, we decompose the effect of $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. Specifically, starting from the raw log probability $\log p(x_{t}|x_{<t})$ (which degenerates to the design of Min-K%), we gradually incorporate $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ into the score computation, until we reach the final formulation of our Min-K%++. For example, when $\log p(x_{t}|x_{<t})$ and $\sigma_{x_{<t}}$ are included (marked by ✓in Table 3), the token-wise score becomes $\frac{\log p(x_{t}|x_{<t})}{\sigma_{x_{<t}}}$. From the results, we see that calibrating the log probability with either $\mu_{x{<t}}$ or $\sigma_{x{<t}}$ alone already leads to 9.3% and 7.0% performance boosts over using raw log probability (i.e., Min-K%). Combining them together, which leads to the formulation specified by Min-K%++, takes advantage from both factors and results in a larger improvement of 16.8%. Such observation again validates the proposed method.
[…]
figureAblation on $k$. Min-K%++ is fairly robust to the hyperparameter choice.
Table 3: Decomposing the token-wise score of Min-K%++ in Equation 5. The first row corresponds to Min-K%, while the last row corresponds to the final design of our Min-K%++.
$\log p(x_{t}|x_{<t})$ $\mu_{x<t}$ $\sigma_{x<t}$ AUROC
68.0
77.3
75.0
84.8
4.4 An online detection setting
Motivation. Just like image generative models typically have a filter to screen harmful generated contents [27], we believe building a similar mechanism for LLMs to detect memorized generated content in an online fashion would be helpful. Say, for example, that after generating a few sentences, the model for some reason suddenly starts to emit copyrighted or private information that is memorized from the training data. In such case, an effective mitigation would be to use pre-training data detection method internally to identify that a part of the generated text is training data, and then let the model refrain from further generation.
Setup. Current benchmarks fail to simulate the online detection setting, since each whole input is either pure training text or pure non-training text. To address this, we adapt WikiMIA to construct an online version of the benchmark. Concretely, each input text is created by concatenating a training text at the end of a non-training text, closely simulating the representative scenario discussed above. Both the training and non-training text have random length, varying among {32, 64, 128}. In this online setting, the prediction on each part of the input, instead of on the whole input, is of interests. Therefore, we split each input into chunks with a length of 32. Methods will be operating on all chunks, and the performance is evaluated on the chunk level. Essentially, this setup simulates using a non-overlapping sliding window of size 32 to sweep over the texts and detecting whether each part within the sliding window is training data or not. See Appendix B Figure 3 for visual illustration.
Table 4: AUROC results in the online detection setting with LLaMA models.
13B 30B 65B
Loss [37] 58.1 61.4 64.0
Zlib [4] 58.3 61.5 64.0
Min-K[28] 58.4 61.6 64.1
Min-K%++ 68.1 67.4 68.4
Results. Table 4 summarizes the results. Note that the Ref, Lowercase, and Neighbor method are no longer applicable or practical in online setting since they require repeated model inference. We see that Min-K%++ is still the most reliable method for online detection. The numbers are lower than those in the conventional offline setting (Table 1), which is expected because now we can only evaluate $\log p(x_{t}|x_{<t},y)$ with $y$ being the prepended text, rather than the exact $\log p(x_{t}|x_{<t})$.
5 Related Work
Membership Inference Attacks. MIA has long been a security and privacy-related topic. Given a target model and a (arbitrary) target input, MIA aims to infer whether the input has been used to train the target model [29, 37]. It has been extensively studied in both vision [5, 6, 8, 38] and language domains [4, 34, 21]. Besides investigation on advanced methodologies, MIA also has profound implications in quantifying privacy risks [23], measuring memorization [7], helping with privacy auditing [30, 36], and detecting test-set contamination [26] and copyrighted contents [22, 11].
Pre-training data detection for LLMs. Despite being an instance of MIA (the general definition and evaluation metrics remain the same), this problem poses unique challenges compared with conventional MIA settings [28, 10]. First, certain assumptions such as no access to subsets of known training data (as nowadays many LLMs’ pre-training corpus is nonpublic) make some early methods that need to train shadow models on the same distribution of the target model impractical. Second, characteristics of LLM pre-training (e.g., few training epochs, large-scale training) inherently makes MIA much more challenging. Lastly, most existing works on MIA against LLMs target the fine-tuning stage [24, 12, 21], which cannot transfer to the pre-training stage for the same reasons.
As a result, there has been growing interests in pre-training data detection recently, though it is still largely underexplored. Shi et al. [28], to our knowledge, is the first to investigate this problem. They contribute the WikiMIA benchmark and propose the Min-Kmethod as a simple yet powerful baseline. Duan et al. [10] construct the MIMIR benchmark, perform systematic evaluation of prior approaches, and analyze the challenges of pre-training data detection for LLMs. In this work we propose a novel method that outperforms Min-Kand achieve superior performances over existing reference-free methods on both benchmarks. There are also works that specifically investigate copy-righted content detection, e.g., identifying memorized books [11].
6 Conclusion and Discussion
In this work, we propose Min-K%++ as a novel method for pre-training data detection for LLMs. Motivated by our insight that maximum likelihood training implicitly optimizes the Hessian trace of likelihood over training data, we design our method to measure how sharply peaked the likelihood is around an input under the discrete distribution of the LLM. It consistently achieves superior performances on two existing benchmarks and in various settings, which establishes a solid baseline for future studies. We hope that our method, along with our theoretical and empirical analysis, can motivate more research upon the pre-training data detection problem.
Limitation. A minor limitation of Min-K%++ is that it cannot be applied to a pure black-box model or API where logits are not available (e.g., OpenAI API only returns at most top-4 token probabilities). However, we emphasize that this not a technical limitation since our method follows the standard grey-box access assumption [21, 28, 10] and is applicable to all open-source LLMs.
Future work. Even with the proposed Min-K%++, detection on MIMIR is still challenging given the limited distribution shifts between training and non-training data. Understanding why and how exactly distribution shifts affect detection remains an important direction for future study. Meanwhile, our theoretical insight discussed in Section 3.2 is around general maximum likelihood training and has strong potential to generalize to other modalities, e.g., generative image modeling [27]. Investigating whether the concept of our method can work for other applications would also be meaningful.
Appendix
Appendix A Further Discussion on Min-K%++
Here we discuss that the numerator of Min-K%++’s formulation in Equation 5 shares a similar form to an approximation of the negative Hessian trace. This further justifies our design since we have shown that the Hessian trace is implicitly optimized over training data during maximum likelihood training.
Proposition 1.
The negative Hessian trace $-\text{tr}(H_{f}(x))$ of a function $f$ can be approximated with $f(x)-\mathbb{E}_{n}f(x+n)$, with mild constraints placed on $n$.
1 is proved by Mitchell et al. [25], and we refer readers to their work for details. If we denote the log likelihood function to be $f$, i.e., $\log p(\cdot|x_{<t})=f(\cdot)$, then the numerator of Min-K%++ token score, $\log p(x_{t}|x_{<t})-\mathbb{E}_{z\sim p(\cdot|x_{<t})}[\log p(z|x_{<t})]$, can be written as $f(x_{t})-\mathbb{E}_{z}f(z)$, where $x_{t}$ is the $t$-th token of the input, and $z$ enumerates all tokens in the vocabulary. This follows a similar form to $f(x)-\mathbb{E}_{n}f(x+n)$, meaning that (in a loose sense) the numerator in Equation 5 is approximating the negative Hessian trace, a statistic that is optimized over training instances and thus can be informative for training data detection.
Appendix B Additional Results
We present additional results here. Tables 5 and 6 show the AUROC results on WikiMIA with all the 10 models in the original and paraphrased setting, respectively. Tables 7 and 8 show the TPR at low FPR results. Table 9 shows the TPR at low FPR results on MIMIR. Lastly, Figure 3 shows a demo of the online detection setting that we have discussed in Section 4.4.
Table 5: Full AUROC results on WikiMIA in the original (verbatim training text) setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 61.0 62.2 60.9 61.9 64.1 63.2 66.8
Mamba-2.8B 64.1 67.0 63.6 64.7 67.0 66.1 69.3
Pythia-2.8B 61.4 61.3 60.9 62.1 64.2 61.8 64.4
Pythia-6.9B 63.8 63.6 62.2 64.3 65.8 66.3 70.3
Pythia-12B 65.4 65.1 64.8 65.8 66.6 68.1 72.3
NeoX-20B 68.8 67.2 68.0 69.0 70.2 71.8 75.0
LLaMA-13B 67.5 57.9 64.0 67.8 65.8 68.0 84.8
LLaMA-30B 69.4 63.5 64.1 69.8 67.6 70.1 84.3
LLaMA-65B 70.7 68.8 66.5 71.1 69.6 71.3 85.1
OPT-66B 65.7 68.7 63.0 66.0 68.2 67.7 70.2
[…]
Average 65.8 64.5 63.8 66.3 66.9 67.4 74.3
64 Mamba-1.4B 58.2 60.6 57.0 60.4 60.6 62.2 67.2
Mamba-2.8B 61.2 64.3 61.7 63.0 63.6 65.4 70.6
Pythia-2.8B 58.4 59.6 57.8 60.6 61.3 61.2 65.0
Pythia-6.9B 60.7 62.4 58.2 62.6 63.2 65.0 71.6
Pythia-12B 61.9 63.0 59.6 63.5 62.6 67.8 72.6
NeoX-20B 66.2 65.7 65.8 67.6 67.1 72.2 76.0
LLaMA-13B 63.6 63.4 62.0 65.3 64.1 66.0 85.7
LLaMA-30B 66.2 69.0 62.1 67.5 67.1 68.5 84.7
LLaMA-65B 67.9 73.4 64.5 69.1 69.6 69.8 83.8
OPT-66B 62.3 67.0 61.2 63.9 64.1 67.0 70.0
[…]
Average 62.7 64.8 61.0 64.4 64.3 66.5 74.7
128 Mamba-1.4B 63.3 62.0 58.5 65.6 64.8 66.8 68.8
Mamba-2.8B 66.2 66.9 62.4 68.5 67.7 71.0 73.4
Pythia-2.8B 62.8 59.6 59.5 65.0 65.2 66.8 66.8
Pythia-6.9B 65.1 63.3 60.5 67.6 67.5 69.5 70.7
Pythia-12B 65.8 63.9 61.4 67.8 67.1 70.7 72.7
NeoX-20B 70.1 67.8 67.7 71.8 71.6 75.0 75.9
LLaMA-13B 67.8 62.6 60.6 69.7 68.3 71.5 83.9
LLaMA-30B 70.3 71.9 59.1 71.8 72.2 73.9 82.6
LLaMA-65B 70.7 73.7 63.3 72.1 73.7 73.6 80.0
OPT-66B 65.5 66.9 59.3 67.5 67.7 70.5 72.3
[…]
Average 66.8 65.8 61.2 68.7 68.6 70.9 74.7
Table 6: Full AUROC results on WikiMIA in the paraphrased setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 61.4 62.3 60.6 62.3 63.6 62.9 66.1
Mamba-2.8B 64.5 66.6 63.5 64.8 66.3 65.3 67.9
Pythia-2.8B 61.6 61.2 60.3 62.3 64.5 61.7 62.4
Pythia-6.9B 64.1 63.5 61.7 64.2 65.5 65.2 68.0
Pythia-12B 65.6 64.9 64.4 65.9 66.8 67.2 69.8
NeoX-20B 68.2 66.3 66.7 68.2 68.3 69.7 69.6
LLaMA-13B 68.0 56.2 63.2 68.3 65.0 68.4 82.7
LLaMA-30B 70.2 62.4 61.2 70.4 66.3 70.7 81.2
LLaMA-65B 71.8 68.2 64.8 72.0 68.7 72.2 81.4
OPT-66B 65.3 68.2 62.7 65.4 66.7 66.3 68.1
[…]
Average 66.1 64.0 62.9 66.4 66.2 67.0 71.7
64 Mamba-1.4B 56.4 59.6 57.0 59.1 60.6 58.0 63.3
Mamba-2.8B 59.8 64.5 62.0 61.9 63.7 62.4 65.8
Pythia-2.8B 56.5 59.2 56.1 59.0 59.6 56.8 58.5
Pythia-6.9B 59.3 62.9 57.7 61.6 63.1 61.1 64.8
Pythia-12B 60.0 63.2 59.1 62.1 62.8 62.5 65.8
NeoX-20B 64.4 65.9 65.1 66.4 67.4 66.1 67.5
LLaMA-13B 63.1 60.9 61.0 65.3 64.7 64.0 78.8
LLaMA-30B 65.5 65.4 59.8 67.4 66.7 65.7 74.9
LLaMA-65B 67.7 71.0 61.9 69.3 69.5 67.9 74.0
OPT-66B 60.4 67.9 60.1 62.3 64.6 62.6 64.7
[…]
Average 61.3 64.0 60.0 63.4 64.3 62.7 67.8
128 Mamba-1.4B 62.7 61.1 57.7 65.3 62.6 64.5 65.6
Mamba-2.8B 65.7 66.6 61.2 68.3 64.6 68.0 70.0
Pythia-2.8B 62.3 59.5 59.6 65.0 61.9 64.7 63.4
Pythia-6.9B 64.7 62.9 60.0 67.4 64.3 67.0 66.8
Pythia-12B 65.4 63.9 60.4 67.9 64.3 68.5 68.8
NeoX-20B 69.5 67.8 67.4 71.8 69.6 72.6 72.2
LLaMA-13B 67.2 59.7 56.4 69.6 64.0 68.7 76.2
LLaMA-30B 69.2 70.0 55.4 71.5 67.2 70.2 73.8
LLaMA-65B 70.2 72.0 60.1 72.1 70.3 70.8 70.7
OPT-66B 64.5 66.8 57.4 66.9 63.4 67.2 68.2
[…]
Average 66.1 65.0 59.5 68.6 65.2 68.2 69.6
Table 7: Full TPR at low FPR (FPR=5%) results on WikiMIA in the original (verbatim training text) setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 14.2 7.8 11.1 15.5 11.9 14.7 12.9
Mamba-2.8B 14.7 9.8 16.8 16.3 16.0 18.1 13.4
Pythia-2.8B 14.7 6.2 11.1 15.8 15.0 17.1 14.2
Pythia-6.9B 14.2 6.7 10.6 16.3 16.5 17.8 17.1
Pythia-12B 17.1 9.0 16.3 17.1 19.4 23.0 18.6
NeoX-20B 19.9 15.5 18.1 19.9 22.2 27.9 19.4
LLaMA-13B 13.9 4.7 9.6 11.6 11.6 18.9 38.5
LLaMA-30B 18.4 9.8 11.4 14.5 9.3 21.2 31.3
LLaMA-65B 22.5 12.4 12.1 18.6 6.5 26.1 41.1
OPT-66B 14.2 10.8 10.6 16.0 21.7 22.0 19.4
[…]
Average 16.4 9.3 12.8 16.1 15.0 20.7 22.6
64 Mamba-1.4B 9.5 4.6 8.8 14.1 8.8 19.4 16.6
Mamba-2.8B 10.2 9.2 16.6 14.8 10.6 19.0 21.5
Pythia-2.8B 10.2 10.6 10.2 14.4 10.2 18.3 16.2
Pythia-6.9B 13.4 12.0 11.6 16.2 10.9 19.0 26.1
Pythia-12B 9.2 13.0 12.3 11.3 11.3 21.5 20.1
NeoX-20B 13.0 15.5 15.5 16.6 13.0 20.4 20.4
LLaMA-13B 11.3 4.2 11.6 12.7 10.2 17.2 34.1
LLaMA-30B 13.7 11.3 11.3 15.5 9.9 17.6 36.3
LLaMA-65B 15.1 13.0 12.3 16.9 9.9 18.0 38.4
OPT-66B 13.4 13.0 10.9 13.4 12.0 26.4 22.5
[…]
Average 11.9 10.6 12.1 14.6 10.7 19.7 25.2
128 Mamba-1.4B 11.5 10.1 12.9 19.4 15.8 16.6 16.6
Mamba-2.8B 19.4 10.1 13.7 23.7 15.1 25.9 21.6
Pythia-2.8B 9.3 10.1 10.8 18.7 8.6 13.7 17.3
Pythia-6.9B 14.4 13.7 12.9 20.9 10.8 18.0 22.3
Pythia-12B 18.0 12.2 12.9 23.7 10.1 25.2 20.9
NeoX-20B 18.7 15.8 12.2 23.0 15.8 25.2 23.0
LLaMA-13B 21.6 10.8 15.8 18.7 12.9 25.9 43.2
LLaMA-30B 23.7 10.8 10.1 18.0 15.1 23.7 40.3
LLaMA-65B 23.0 18.0 14.4 22.3 15.8 23.7 27.3
OPT-66B 20.9 17.3 14.4 21.6 12.9 23.0 16.6
[…]
Average 18.1 12.9 13.0 21.0 13.3 22.1 24.9
Table 8: Full TPR at low FPR (FPR=5%) results on WikiMIA in the paraphrased setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 14.2 5.9 13.2 13.2 7.2 15.2 10.6
Mamba-2.8B 16.5 10.1 15.0 12.7 9.3 19.9 13.4
Pythia-2.8B 14.2 7.2 11.6 14.5 8.5 16.5 13.9
Pythia-6.9B 15.0 6.2 11.9 12.7 9.6 21.7 17.1
Pythia-12B 17.3 8.0 16.5 15.5 9.8 19.9 17.3
NeoX-20B 18.1 15.2 15.5 18.6 15.2 19.6 12.9
LLaMA-13B 16.3 5.4 9.6 15.0 8.5 17.6 35.9
LLaMA-30B 14.7 7.5 12.7 15.0 9.3 18.1 27.4
LLaMA-65B 23.3 9.3 11.9 16.5 12.1 24.3 35.9
OPT-66B 15.2 10.3 13.4 17.1 12.1 18.1 15.2
[…]
Average 16.5 8.5 13.1 15.1 10.2 19.1 20.0
64 Mamba-1.4B 8.1 8.1 9.5 15.1 9.5 8.4 7.0
Mamba-2.8B 12.3 11.3 14.8 14.8 18.3 13.0 12.3
Pythia-2.8B 9.5 13.0 11.3 16.6 11.3 11.3 9.9
Pythia-6.9B 10.6 16.2 11.3 15.8 12.7 12.7 14.1
Pythia-12B 11.6 14.4 13.4 16.2 10.6 14.4 13.7
NeoX-20B 16.2 14.1 13.7 19.4 18.3 17.6 13.0
LLaMA-13B 12.0 4.6 13.7 13.4 14.4 13.4 26.4
LLaMA-30B 13.4 8.1 8.1 16.9 11.6 14.4 21.5
LLaMA-65B 13.4 10.9 9.5 18.0 16.9 13.7 29.2
OPT-66B 13.4 13.0 13.4 14.8 13.7 14.8 12.7
[…]
Average 12.0 11.4 11.9 16.1 13.7 13.4 16.0
128 Mamba-1.4B 13.7 11.5 11.5 17.3 13.7 14.4 10.1
Mamba-2.8B 16.6 10.8 15.1 20.1 17.3 20.1 15.1
Pythia-2.8B 14.4 7.2 8.6 16.6 12.2 14.4 14.4
Pythia-6.9B 16.6 8.6 11.5 20.9 17.3 17.3 21.6
Pythia-12B 19.4 8.6 12.2 19.4 10.1 21.6 17.3
NeoX-20B 15.8 19.4 16.6 21.6 18.7 22.3 19.4
LLaMA-13B 18.0 4.3 15.8 21.6 13.7 20.1 35.2
LLaMA-30B 18.7 18.7 13.7 19.4 14.4 18.7 21.6
LLaMA-65B 24.5 12.9 13.7 22.3 18.7 25.2 25.2
OPT-66B 18.0 15.8 11.5 18.7 12.9 20.1 18.7
[…]
Average 17.6 11.8 13.0 19.8 14.9 19.4 19.9
Table 9: TPR at low FPR (FPR=5%) results on MIMIR. In each column, the best result across all methods is bolded, with the runner-up underlined. Ref relies on a reference LLM. Neighbor induces significant extra computational cost than others (25×25$\times$25 × in this case), for which reason we don’t run on the 12B model.
Wikipedia Github Pile CC PubMed Central
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [37] 4.2 4.7 4.7 5.1 5.0 22.6 32.1 33.6 38.5 40.7 3.1 5.0 4.8 4.9 5.1 4.0 4.4 4.3 4.9 5.0
Ref [4] 6.1 5.3 5.5 5.6 5.7 23.4 14.8 14.9 15.4 16.2 5.5 5.6 5.8 5.8 7.5 5.7 4.1 4.0 5.9 8.7
Zlib [4] 4.2 5.7 5.9 6.3 6.8 25.0 32.8 36.1 39.3 40.8 4.0 5.1 5.4 6.2 6.6 3.8 3.6 3.5 4.3 4.4
Neighbor [21] 4.0 4.5 4.9 5.8 / 24.7 31.6 29.8 34.1 / 3.9 3.6 4.0 5.3 / 3.9 3.7 4.5 4.5 /
Min-K[28] 6.4 5.6 6.4 6.5 8.1 23.3 32.2 34.0 39.0 40.8 4.2 5.1 5.2 5.5 5.7 4.7 5.2 4.8 5.9 5.4
Min-K%++ 5.7 6.1 8.5 11.4 11.5 25.4 33.2 34.2 38.2 40.1 5.8 5.0 5.4 6.3 6.3 5.1 6.3 6.5 7.4 9.0
ArXiv DM Mathematics HackerNews Average
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [37] 4.0 4.8 4.6 5.4 5.6 3.8 4.3 4.1 4.1 4.0 5.0 4.8 5.5 5.9 6.8 6.7 8.6 8.8 9.8 10.3
Ref [4] 5.2 5.4 5.9 6.8 7.4 5.3 3.6 4.9 5.4 5.9 5.2 6.3 7.6 7.0 7.6 8.1 6.4 6.9 7.4 8.4
Zlib [4] 2.9 4.3 4.1 4.6 4.7 4.1 5.0 4.6 4.3 4.3 5.0 5.5 5.8 5.6 5.8 7.0 8.9 9.3 10.1 10.5
Neighbor [21] 4.7 4.8 4.4 4.1 / 5.6 4.4 4.5 4.5 / 6.5 5.2 5.3 5.7 / 7.6 8.3 8.2 9.1 /
Min-K[28] 4.9 4.8 4.7 5.6 6.2 4.5 4.5 4.6 4.7 5.2 5.2 5.7 5.9 6.3 6.9 7.6 9.0 9.4 10.5 11.2
Min-K%++ 6.0 6.0 6.7 8.2 8.6 5.4 5.5 5.7 6.2 6.3 5.5 4.9 5.7 6.6 6.6 8.4 9.6 10.4 12.0 12.6
Figure 3: A demo of Min-K%++ in the online detection setting. Each input is split into 32-length chunks or sliding windows (separated with the mark ‘|’), and Min-K%++ performs detection over each chunk. The token-wise scores that contribute to the chunk-wise score are marked by the darker colors. The detection results are marked by non-training data v.s. training data. During LLM’s generation, once we detect training data in the current chunk we can stop further generation to avoid potential issues.