### v3

# Min-K%++: Improved Baseline for Detecting Pre-Training Data from Large Language Models

Abstract
The problem of pre-training data detection for large language models (LLMs) has received growing attention due to its implications in critical issues like copyright violation and test data contamination. Despite improved performance, existing methods (including the state-of-the-art, Min-K%) are mostly developed upon simple heuristics and lack solid, reasonable foundations. In this work, we propose a novel and theoretically motivated methodology for pre-training data detection, named Min-K%++. Specifically, we present a key insight that training samples tend to be local maxima of the modeled distribution along each input dimension through maximum likelihood training, which in turn allow us to insightfully translate the problem into identification of local maxima. Then, we design our method accordingly that works under the discrete distribution modeled by LLMs, whose core idea is to determine whether the input forms a mode or has relatively high probability under the conditional categorical distribution. Empirically, the proposed method achieves new SOTA performance across multiple settings. On the WikiMIA benchmark, Min-K%++ outperforms the runner-up by 6.2% to 10.5% in detection AUROC averaged over five models. On the more challenging MIMIR benchmark, it consistently improves upon reference-free methods while performing on par with reference-based method that requires an extra reference model.
1 Introduction
Data is one of the most important factors for the success of large language models (LLMs). As the training corpus grows in scale, it has increasing tendency to be held in-house as proprietary data instead of being publicly disclosed [33, 1]. However, for large-scale training corpora that consist of up to trillions of tokens [9], the sheer volume of the training corpus can lead to unintended negative consequences. For example, memorized private information is vulnerable to data extraction [4], and memorized copyrighted contents (e.g., books and news articles) may violate the rights of content creators [14, 20]. Furthermore, it becomes increasingly likely that evaluation data is exposed at training time, bringing the faithfulness and effectiveness of evaluation benchmarks into question [27].
For these reasons, there has been growing interest in effective pre-training data detection strategies. Pre-training data detection can be considered a special case of Membership Inference Attack (MIA) [30]: the goal is to infer whether a given input has been used for training a target LLM (see Figure 1 left for illustration). Due to characteristics of pre-training corpora and training characteristics of LLMs [29, 10], this problem has been shown to be much more challenging than conventional MIA settings (see Section 5 for details). There are a few methods proposed recently, dedicated to this problem [4, 23, 29]. However, despite their improved performance, most existing methods are developed with simple heuristics, lacking solid and interpretable foundations.
In this work, we propose a novel and theoretically motivated methodology, Min-K%++, for pre-training data detection. Our exploration starts by asking a fundamental question of “what characteristics or footprints do training samples exhibit within the model”, where we answer it by revisiting the maximum likelihood training objective through the lens of score matching [18]. Importantly, we uncover that for continuous distributions with maximum likelihood training, training data points tend to be local maxima or be located near local maxima along input dimensions, a key insight that allows us for the first time to translate the training data detection problem into identification of local maxima. Then, upon theoretical insight, we develop a practical scoring mechanism that works for the discrete distribution modeled by LLMs. The core idea is to examine whether the input next token forms a mode or has relatively high probability under the conditional categorical distribution.
Figure 1: Left: We study the pre-training data detection problem for LLMs [29, 10]. Given the grey-box access to the target LLM (only the logits, token probabilities, and loss are available), the goal is to design a scoring function that yields a score which (upon thresholding) best separates training and non-training text. Right: ROC (receiver operating characteristic) curves of various methods on WikiMIA benchmark [29]. The AUROC (area under the ROC curve) score is denoted in the legend. Our Min-K%++ improves upon existing approaches by large margin.
Empirically, we perform extensive experiments to verify the proposed method on two established benchmarks with up to 5 families of models, including LLaMA [32], Pythia [2], and the new architecture Mamba [15]. Figure 1 right showcases the performance comparison between our method and previous ones. Notably, Min-K%++ on average leads to 6.2% to 10.5% absolute increases over the runner-up Min-K[29] on the WikiMIA benchmark [29]. On the more challenging MIMIR benchmark, our method still consistently outperforms other reference-free methods and is on par with the reference-based method [4]: unlike reference-based methods which require another LLM to calibrate the likelihood, Min-K%++ is applied to a standalone target LLM. We also introduce and study an online detection setting that simulates “detect-while-generating”; the proposed Min-K%++ again performs the best. Lastly, Min-K%++ can also be interpreted from a calibration perspective: with ablation study we show that both calibration factors in the formulation contribute to the overall high performance. We summarize our contributions as follows:
1) We establish that for maximum likelihood training, training instances tend to form local maxima or locate near local maxima along each input dimension (under continuous distributions), allowing us to turn training data detection into local maxima identification. 2) We develop a novel and sound methodology for LLM training data detection which functions by examining whether the input forms a mode or has relatively high probability under the conditional categorical distribution. 3) We demonstrate remarkable improvements over existing methods on established benchmarks.
2 Background
In this section, we first cover the problem statement of pre-training data detection defined by prior works [30, 29, 10]. Then we briefly introduce how LLMs work and how they are trained.
Problem statement. Pre-training data detection is an instance of Membership Inference Attack (MIA) [30]. Formally, given 1) a data instance $x$ and 2) a pre-trained auto-regressive LLM $\mathcal{M}$ that is trained on a dataset $\mathcal{D}$, the goal is to infer whether $x\in\mathcal{D}$ or not (i.e., $x$ is training data or non-training data). The approach to detection leverages a scoring function $s(x;\mathcal{M})$ that computes a score for each input. A threshold is then applied to the score to yield a binary prediction:
$\text{prediction}(x,\mathcal{M})=\begin{cases}1~{}(x\in\mathcal{D}),&s(x;\mathcal{M})\geq\lambda\\ 0~{}(x\notin\mathcal{D}),&s(x;\mathcal{M})<\lambda\end{cases},$ (1)
where $\lambda$ is a case-dependent threshold. Following the established standard [23, 29, 10], we consider grey-box access to the target model $\mathcal{M}$, meaning that one can only access the output statistics including the loss value, logits, and token probabilities. Additional information such as the model weights and gradients are not available. In summary, the key to the pre-training data detection is to design an appropriate scoring function that best separates training data from non-training data.
(Auto-regressive) LLMs. LLMs are typically trained by maximum likelihood training that maximizes the probability of training token sequences [35]. In particular, auto-regressive LLMs decompose the probability of token sequence $(x_{1},x_{2},...,x_{t})$ with chain rule [37], i.e., $p(x_{1},x_{2},...,x_{t})=p(x_{t}|x_{1},x_{2},...x_{t-1})\cdot p(x_{1},x_{2},...x_{t-1})$. For brevity, throughout the paper we abbreviate the prefix of $x_{t}$ as $x_{<t}$. At inference stage, LLMs yield new tokens one by one according to its predicted conditional categorical distribution $p(\cdot|x_{<t})$ over the vocabulary.
3 Min-K%++
In this section we introduce our methodology, called Min-K%++. We start by discussing our grounding observation around general maximum likelihood training for continuous distribution, where we unveil that training samples tend to form local maxima or locate near local maxima along each input dimension of the distribution captured by the model. This insight allows us to introduce a unique and theoretically sound perspective for addressing the problem. Then, translating our insight into a practical solution, we formulate the proposed Min-K%++ which works for discrete inputs of LLMs by examining whether the input forms a mode or has a relatively high probability under the conditional categorical distribution of the predicted next token.
3.1 Motivation & Insight
To approach the pre-training data detection problem, unlike prior methods that often rely on heuristics [39, 5, 29], we choose to start our in-depth exploration by asking a fundamental question that existing works fail to touch to our knowledge. That is, what characteristics or “footprints” do training samples leave to the model after training? Since training data interacts with the model through training-time optimization, to answer the question, we revisit the maximum likelihood training objective of (auto-regressive) LLMs through the lens of score matching [18]. For the ease of discussion, let us generalize from LLMs, which model discrete distribution over tokens, to the general maximum likelihood training with continuous distribution for now.
Earlier, Koehler et al. [21] have proved that implicit score matching (ISM) objective [18] is a relaxation (within a multiplicative factor) of maximum likelihood estimation. Consequently, one can reformulate the maximum likelihood training loss with ISM as
$\frac{1}{N}\sum_{\bm{x}}\biggl{[}\frac{1}{2}||\psi(\bm{x})||^{2}+\underbrace{\sum_{i=1}^{d}\frac{\partial\psi_{i}(\bm{x})}{\partial\bm{x}_{i}}}_{\mathclap{\begin{subarray}{c}\text{the sum of the second-order}\\ \text{partial derivatives}\end{subarray}}}\biggr{]},$ (2)
where $\psi(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}}$ is the score function [18] for a $d$-dimensional input $\bm{x}$, $\bm{x}_{i}$ is the element at the $i$-th dimension of $\bm{x}$, $\psi_{i}(\bm{x})=\frac{\partial\log p(\bm{x})}{\partial\bm{x}_{i}}$, and $N$ is the number of training samples. Without ambiguity, we are omitting the model parameters in the formulation of $\psi$ for brevity.
Remark. It can be seen from Equation 2 that maximum likelihood training implicitly minimizes both 1) the magnitude of first-order derivatives of likelihood $\log p(\bm{x})$ w.r.t. $\bm{x}$ and 2) the sum of the second-order partial derivatives of $\log p(\bm{x})$ w.r.t. each dimension of $\bm{x}$. As a result, we posit that the first-order $\frac{\partial\log p(\bm{x})}{\partial\bm{x}_{i}}$ will be close to 0 and the second-order $\frac{\partial^{2}\log p(\bm{x})}{\partial\bm{x}_{i}^{2}}$ will be minimized to be negative for training samples. Importantly, according to the univariate second-derivative test [34], the implication here is that when looking along each dimension of the input (corresponding to each $\bm{x}_{i}$), training data points tend to form local maxima or be located near local maxima of the likelihood. Given this key insight, we can translate the problem of detecting training data into identifying whether the input is a local maximum or locates near a local maximum (along input dimensions), providing a novel and theoretically grounded perspective for tackling this task. Note that although we focus on LLMs in this work, our discussion here is general and can be applied to other modalities and models, e.g., VAEs [19] and diffusion models [17, 28] for image modeling, which are trained by optimizing the evidence lower bound of likelihood [19, 17].
3.2 Formulating Min-K%++
From insights to an actual method. In the context of LLMs, the dimension of a token sequence corresponds to each token. Also, note that tokens are discrete rather than continuous and there is no notion of locality (local maxima) anymore. Therefore, to determine whether a token sequence $(x_{<t},x_{t})$ has appeared in training, we choose to examine whether it forms a mode or has a relatively high probability along the dimension that corresponds to $x_{t}$. Again, in our context, moving along the dimension of a token means replacing the current one with other candidate tokens within the vocabulary of the model. Our idea here closely reflects the insight in the continuous distribution case we have discussed above, where we proposed to detect training data by identifying if the input is a local maximum or locates near a local maximum.
Meanwhile, since the probability $p(x_{<t},x_{t})$ can be decomposed by $p(x_{<t},x_{t})=p(x_{t}|x_{<t})\cdot p(x_{<t})$, the problem further transforms into identifying if the next token $x_{t}$ forms a mode or has a relatively high probability under the conditional categorical distribution $p(\cdot|x_{<t})$. To this end, we propose the following formulation for our Min-K%++:
[…]
Here, $\mu_{x_{<t}}=\mathbb{E}_{z\sim p(\cdot|x_{<t})}[\log p(z|x_{<t})]$ is the expectation of the next token’s log probability over the vocabulary of the model given the prefix $x_{<t}$, and $\sigma_{x_{<t}}=\sqrt{\mathbb{E}_{z\sim p(\cdot|x_{<t})}[(\log p(z|x_{<t})-\mu_{x_{<t}})^{2}]}$ is the standard deviation. In practice, both terms can be computed analytically since the categorical distribution $p(\cdot|x_{<t})$ is encoded by the output logits of the model, which we have access to. Upon the sequence-wise score computed by Equation 3, we adopt a similar strategy to Shi et al. [29] where we select the $k$of the token sequences with the minimum score and compute the average over them to obtain the final sentence-wise score, as described by Equation 4. This is because in practice LLM training minimizes the loss of all possible sequences in each training text/sentence. Therefore, by computing an aggregated score we expect to get a more robust estimation for the full sentence input. Our method is closely motivated by our theoretical insight. We provide further interpretations for more comprehensive understanding.
Interpretation 1: More robust identification of the mode. We have translated training data detection into determining whether the input next token forms a mode of the conditional distribution. Besides motivating our method, this unique scheme also allows us to understand why existing methods may be sub-optimal (which was unclear earlier since they were mostly developed upon heuristics).
Specifically, let us take the state-of-the-art Min-K[29] as an example, which follows a simple intuition that training texts are less likely to include low-probability “outlier” tokens and thus uses the token probability as the score. Within our proposed scheme, this may not identify the mode robustly or accurately as token probabilities can vary a lot across different inputs. For example, some training texts may be rare or inherently difficult to learn, thus having low token probabilities despite still being the mode. Our design, in contrast, can account for such cases since it explicitly compares the log likelihood of the target next token $\log p(x_{t}|x_{<t})$ and the expected log likelihood over all candidate tokens $\mu_{x_{<t}}$. Regardless of the absolute value of the token likelihood, the larger $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$ is, the more likely $x_{t}$ forms a mode or has a relatively high probability among the vocabulary.
To see this more clearly, let us consider a conceptual example depicted in Figure 2, where we show the predicted next token probability distribution for two different input token sequences $(x_{<t},x_{t})$, supposing a vocabulary of size 10. In the two cases (Figure 2 (a) and (b)), both $p(x_{t}|x_{<t})$ is 0.2. Existing methods often measure exactly the next token probability [29] and thus will give the same judgement to both inputs (e.g., that the sequence $(x_{<t},x_{t})$ was not seen in the training). However, inspecting them with the idea of our Min-K%++ reveals their distinct characteristics:
  • In Figure 2 (a), one can notice that the distribution is actually peaked at $x_{t}$ (the bar with the hatch pattern ‘//’), indicating that the current sequence $x_{t}$ forms a mode of $p(\cdot|x_{<t})$ and is very likely to have appeared in the training.
  • In Figure 2 (b), the probability distribution is not peaked at $x_{t}$ as there is another token that gets assigned much larger probability mass. This means that $x_{t}$ is less likely to follow the prefix $x_{<t}$ according to the learned distribution of the model, which should be seen as an indicator that the sequence $(x_{<t},x_{t})$ was not seen in the training.
The above reasoning can be concisely reflected by comparing $\log p(x_{t}|x_{<t})$ and $\mu_{x_{<t}}$, which are marked by the two dashed lines in Figure 2. Accordingly, it can be seen that Figure 2 (a) and (b) exhibit positive and negative $\log p(x_{t}|x_{<t})-\mu_{x_{<t}}$, respectively, which successfully separate them apart.
Interpretation 2: A calibration perspective. Reference-based membership inference attacks [4] show superior performance by calibrating the (sentence-level) likelihood with certain “references”, such as the likelihood of the same input on a smaller LLM, the Zlib entropy, or the likelihood of lowercased text. In this regard, Min-K%++ can be seen as a calibrating method, too, where it calibrates the next token log likelihood with two calibration factors, $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. The unique advantage of our method lies in that instead of using some external references (e.g., an extra LLM), we leverage the statistics that can be readily computed with the target model only. The effect of $\mu_{x_{<t}}$ has been thoroughly discussed in the first interpretation. The inclusion of $\sigma_{x_{<t}}$ is inspired by temperature scaling [16], a technique for calibrating the prediction confidence of neural networks. It scales the model output by a constant and has been shown to benefit tasks like out-of-distribution detection [22, 41]. Instead of using a constant, here we use a dynamic factor $\sigma_{x_{<t}}$ that is adaptive to each input and find that it leads to extra performance boosts. Later we will demonstrate that both factors contribute to the success of Min-K%++ through ablation study.
Figure 2: A conceptual example showcasing the idea of Min-K%++. (a) and (b) show two next token distributions for two different input token sequences $(x_{<t},x_{t})$. The bars with the hatch pattern (‘//’) correspond to the target next token $x_{t}$. Existing methods often measure the exact next token probability of $x_{t}$ (hatched bars), which fails to separate these two inputs due to the same probability (0.2). By comparison, Min-K%++ determines if $x_{t}$ forms a mode or has relatively high probability by comparing $\log p(x_{t}|x_{<t})$ with $\mu_{x_{<t}}$; the score is the difference between the red and pink dashed lines. See text for detailed discussion.
4 Experiments
We conduct extensive experiments to validate the empirical efficacy of the proposed Min-K%++. We first focus on two established benchmarks for pre-training data detection. Following that, ablation studies are presented and discussed. Lastly, we investigate a new online detection setting.
4.1 Setup
Benchmarks. We focus on two benchmarks (and the only two to our knowledge) for pre-training data detection, WikiMIA [29] and MIMIR [10]. WikiMIA is the first benchmark for pre-training data detection, which consists of texts from Wikipedia events. The training v.s. non-training data is determined by the timestamp. WikiMIA specifically groups data into splits according to the sentence length, intending to provide a fine-grained evaluation. It also considers two settings: original and paraphrased. The former assesses the detection of verbatim training texts, while the latter paraphrases the training texts (using ChatGPT) and evaluates on paraphrased inputs.
MIMIR [10] is built upon the Pile dataset [13], where training samples and non-training samples are drawn from the train and test split, respectively. MIMIR is found to be much more challenging than WikiMIA since the training and non-training texts are from the same dataset and thus have minimal distribution shifts and temporal discrepancy [10].
Baselines. We consider 6 representative and state-of-the-art methods as our baselines, which are also featured in the benchmarking work of Duan et al. [10]. Except Min-K%, all methods consider sentence-level probability. Loss method [39] is a general technique that directly takes the loss as the score for detection. Reference method [4] (Ref) uses a reference LLM to calibrate the likelihood of the input. Zlib and Lowercase method [4] are also reference-based, except that they use zlib compression entropy and the likelihood of lowercased text as the reference, respectively. Neighbor method [23] perturbs the input sentence with masked language models to create “neighbors” and calibrate the loss of the input sentence with the average loss of the neighbor sentences. Lastly, Min-K[29] examines the exact token probability and averages a subset of minimum token scores over the input; it is currently the best-performing method on WikiMIA.
Determining the $k$%. Our Min-K%++ has a hyperparameter $k$ (Equation 4). As existing benchmarks do not provide validation data, we sweep over 10, 20, …, 100 for $k$ and report the best result. For fair comparison, we also report the best possible performance of other methods. Later we show that Min-K%++ is robust to this hyperparameter.
Models. WikiMIA is applicable to a wide range of models since Wikipedia dumps are often included into the training corpus of many LLMs. Specifically, we consider the suite of Pythia models [2] (2.8B, 6.9B, 12B), GPT-NeoX [3] (20B), LLaMA [32] (13B, 30B, 65B), and OPT [42] (66B). We further include a new family of architectures, Mamba [15]. When a reference model is needed, following Shi et al. [29] we use the smaller version correspondingly, e.g., LLaMA-7B for LLaMA models and Pythia-70M for Pythia models. MIMIR is applicable to models that are trained on Pile. To be consistent with Duan et al. [10], we focus on Pythia models (160M, 1.4B, 2.8B, 6.9B, 12B).
Metrics. As a binary classification problem, the detection performance can be evaluated with the AUROC score (area under the receiver operating characteristic curve) [4, 29, 10]. We define training data as “positive” and non-training data as “negative”. AUROC is threshold-independent and can be interpreted as the the probability that a positive instance has higher score than a negative instance according to the detector. Therefore, the higher the better, and the random-guessing baseline is 50%. While we use AUROC as the main metric, we also report True Positive Rate (TPR) at low False Positive Rate (FPR) which measures detection rate at a meaningful threshold.
Compute. All experiments are run with a single Tesla A100 80 GB GPU. When the model is large, we use either half-precision floating point or 8-bit quantization to fit the model into GPU memory.
Table 1: AUROC results on WikiMIA benchmark [29]. Ori. and Para. denote the original and paraphrased settings, respectively. Bolded number shows the best result within each column across all methods. The proposed Min-K%++ leads to remarkable improvements over existing methods in most settings.
Mamba-1.4B Pythia-6.9B LLaMA-13B LLaMA-30B LLaMA-65B Average
Len. Method Ori. Para. Ori. Para. Ori. Para. Ori. Para. Ori. Para. Ori. Para.
32 Loss [39] 61.0 61.4 63.8 64.1 67.5 68.0 69.4 70.2 70.7 71.8 66.5 67.1
Ref [4] 62.2 62.3 63.6 63.5 57.9 56.2 63.5 62.4 68.8 68.2 63.2 62.5
Lowercase [4] 60.9 60.6 62.2 61.7 64.0 63.2 64.1 61.2 66.5 64.8 63.5 62.3
Zlib [4] 61.9 62.3 64.3 64.2 67.8 68.3 69.8 70.4 71.1 72.0 67.0 67.4
Neighbor [23] 64.1 63.6 65.8 65.5 65.8 65.0 67.6 66.3 69.6 68.7 66.6 65.8
Min-K[29] 63.2 62.9 66.3 65.2 68.0 68.4 70.1 70.7 71.3 72.2 67.8 67.9
Min-K%++ 66.8 66.1 70.3 68.0 84.8 82.7 84.3 81.2 85.1 81.4 78.3 75.9
64 Loss [39] 58.2 56.4 60.7 59.3 63.6 63.1 66.2 65.5 67.9 67.7 63.3 62.4
Ref [4] 60.6 59.6 62.4 62.9 63.4 60.9 69.0 65.4 73.4 71.0 65.8 63.9
Lowercase [4] 57.0 57.0 58.2 57.7 62.0 61.0 62.1 59.8 64.5 61.9 60.8 59.5
Zlib [4] 60.4 59.1 62.6 61.6 65.3 65.3 67.5 67.4 69.1 69.3 65.0 64.5
Neighbor [23] 60.6 60.6 63.2 63.1 64.1 64.7 67.1 66.7 69.6 69.5 64.9 64.9
Min-K[29] 62.2 58.0 65.0 61.1 66.0 64.0 68.5 65.7 69.8 67.9 66.3 63.3
Min-K%++ 67.2 63.3 71.6 64.8 85.7 78.8 84.7 74.9 83.8 74.0 78.6 71.2
128 Loss [39] 63.3 62.7 65.1 64.7 67.8 67.2 70.3 69.2 70.7 70.2 67.4 66.8
Ref [4] 62.0 61.1 63.3 62.9 62.6 59.7 71.9 70.0 73.7 72.0 66.7 65.1
Lowercase [4] 58.5 57.7 60.5 60.0 60.6 56.4 59.1 55.4 63.3 60.1 60.4 57.9
Zlib [4] 65.6 65.3 67.6 67.4 69.7 69.6 71.8 71.5 72.1 72.1 69.4 69.2
Neighbor [23] 64.8 62.6 67.5 64.3 68.3 64.0 72.2 67.2 73.7 70.3 69.3 65.7
Min-K[29] 66.8 64.5 69.5 67.0 71.5 68.7 73.9 70.2 73.6 70.8 71.0 68.2
Min-K%++ 68.8 65.6 70.7 66.8 83.9 76.2 82.6 73.8 80.0 70.7 77.2 70.6
Table 2: AUROC results on the challenging MIMIR benchmark [10] with a suite of Pythia models [2]. In each column, the best result across all methods is bolded, with the runner-up underlined. Ref relies on a reference LLM. Neighbor induces significant extra computational cost than others (25×25$\times$25 × in this case), for which reason we don’t run on the 12B model. Despite not requiring a reference model like the Ref method does, our Min-K%++ consistently achieves superior or comparable performance.
Wikipedia Github Pile CC PubMed Central
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [39] 50.2 51.3 51.8 52.8 53.5 65.7 69.8 71.3 73.0 74.0 49.6 50.0 50.1 50.7 51.1 49.9 49.8 49.9 50.6 51.3
Ref [4] 51.2 55.2 58.1 61.8 63.9 63.9 67.1 65.3 64.4 63.0 49.2 52.2 53.7 54.9 56.7 51.3 53.1 53.7 55.9 58.2
Zlib [4] 51.1 52.0 52.4 53.5 54.3 67.4 71.0 72.3 73.9 74.8 49.6 50.1 50.3 50.8 51.1 49.9 50.0 50.1 50.6 51.2
Neighbor [23] 50.7 51.7 52.2 53.2 / 65.3 69.4 70.5 72.1 / 49.6 50.0 50.1 50.8 / 47.9 49.1 49.7 50.1 /
Min-K[29] 50.2 51.3 51.8 53.6 54.4 65.7 69.9 71.4 73.2 74.3 50.3 51.0 50.8 51.5 51.7 50.6 50.3 50.5 51.2 52.3
Min-K%++ 49.7 53.7 55.1 58.0 61.1 64.8 69.6 70.9 72.8 74.2 50.6 51.0 51.0 53.0 53.5 50.6 51.4 52.4 54.2 55.4
ArXiv DM Mathematics HackerNews Average
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [39] 51.0 51.5 51.9 52.9 53.4 48.8 48.5 48.4 48.5 48.5 49.4 50.5 51.3 52.1 52.8 52.1 53.1 53.5 54.4 54.9
Ref [4] 49.4 51.5 53.1 55.8 57.5 51.1 51.1 50.5 51.1 50.9 49.1 52.2 55.1 57.9 60.6 52.2 54.6 55.6 57.4 58.7
Zlib [4] 50.1 50.9 51.3 52.2 52.7 48.1 48.2 48.0 48.1 48.1 49.7 50.3 50.8 51.2 51.7 52.3 53.2 53.6 54.3 54.8
Neighbor [23] 50.7 51.4 51.8 52.2 / 49.0 47.0 46.8 46.6 / 50.9 51.7 51.5 51.9 / 52.0 52.9 53.2 53.8 /
Min-K[29] 51.0 51.7 52.5 53.6 54.6 49.4 49.7 49.5 49.6 49.7 50.9 51.3 52.6 53.6 54.6 52.6 53.6 54.2 55.2 55.9
Min-K%++ 50.1 51.1 53.7 55.2 58.0 50.5 50.9 51.7 51.6 51.9 50.7 51.3 52.6 54.5 56.5 52.4 54.1 55.3 57.0 58.7
4.2 Results
WikiMIA results. Table 1 shows major results in terms of AUROC; for results on more models and TPR numbers, please see Appendix C Tables 5, 6, 7 and 8. We remark that Min-K%++ achieves significant improvements over existing methods. In the original setting, Min-K%++ on average outperforms the runner-up Min-Kby {10.5%, 12.3%, 6.2%} with inputs of length {32, 64, 128}, respectively. In the paraphrased setting, Min-K%++ is also the best-performing approach on average in all cases.
Min-K%++ is consistent across models and input lengths. It can be noticed that Min-K%++’s superior results are agnostic to models: besides transformer-based LLMs, Min-K%++ also decently outperforms others on the new state space-based architecture, Mamba. In terms of input length, Shi et al. [29] identify that short inputs are more challenging than longer inputs. While this is indeed the case for Min-K%, which yields {62.2%, 66.8%} AUROC on {64, 128}-length inputs with Mamba-1.4B (a 4.6% decrease when changing to shorter inputs), Min-K%++ achieves a much more consistent performances of {67.2%, 68.8%} (a mere decrease of 1.6%). Both observations demonstrate the robustness and generality of our proposed method.
MIMIR results. Table 2 shows the AUROC results; see Appendix C Table 9 for TPR results. Most numbers are taken from those reported by Duan et al. [10]. MIMIR is extremely challenging in that the training and non-training texts are maximally similar to each other since they are drawn from the same dataset. Nonetheless, Min-K%++ still improves upon existing methods in most cases. Averaged over 7 subsets, Min-K%++’s relative AUROC w.r.t. the powerful Min-Kis {–0.2%, +0.5%, +1.1%, +1.8%, +2.8%} on Pythia model with {160M, 1.4B, 2.8B, 6.9B, 12B} parameters, respectively. Extrapolating this trend, we anticipate Min-K%++’s effectiveness to be even more obvious with larger models. The most powerful method on MIMIR is the Reference approach [4] (Ref), the results of which are obtained by exhaustively trying out 8 different LLMs as the reference model and picking the best one [10] (note that inferencing with an extra LLM is expensive and may not always be feasible). In contrast, our Min-K%++ does not rely on a reference model, yet provides competitive performance that is on par with Ref. Meanwhile, we remark that Min-K%++ achieves new SOTA results among all the 5 reference-free methods out of the 6 total baselines. Also, in Appendix C Table 9, Min-K%++ outperforms all baselines when it comes to TPR performance.
4.3 Ablation study
We focus on WikiMIA with LLaMA-13B model for ablation study.
Ablation on the hyperparameter $k$%. $k$ determines what percent of token sequences with minimum scores are chosen to compute the final score. From Table 3, it is obvious that Min-K%++ is robust to the choice of $k$, with the best and the worst result being 84.8% and 82.1% (a variation of 2.7%), respectively. Min-Khas a similar hyperparameter but is more sensitive to it: the variation between the best (68.0%) and the worst result (63.6%) is 4.4%, slightly larger than that of Min-K%++.
Decomposing the contribution of calibration factors. Recall that our method can be interpreted as calibrating the log probability $\log p(x_{t}|x_{<t})$ with two calibrationg factors $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ (Equation 3). In Table 3, we decompose the effect of $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$. Specifically, starting from the raw log probability $\log p(x_{t}|x_{<t})$, we gradually incorporate $\mu_{x_{<t}}$ and $\sigma_{x_{<t}}$ into the score computation, until we reach the final formulation of our Min-K%++. For example, when $\log p(x_{t}|x_{<t})$ and $\sigma_{x_{<t}}$ are included (marked by ✓in Table 3), the token-wise score becomes $\frac{\log p(x_{t}|x_{<t})}{\sigma_{x_{<t}}}$. From the results, we see that calibrating the log probability with either $\mu_{x{<t}}$ or $\sigma_{x{<t}}$ alone already leads to 9.3% and 7.0% performance boosts over using raw log probability. Combining them together, which leads to the formulation specified by Min-K%++, takes advantage from both factors and results in a larger improvement of 16.8%. Such observation again validates the proposed method.
[…]
figureAblation on $k$. Min-K%++ is fairly robust to the hyperparameter choice.
Table 3: Decomposing the token-wise score of Min-K%++ in Equation 3. The last row corresponds to the final design of our Min-K%++.
$\log p(x_{t}|x_{<t})$ $\mu_{x<t}$ $\sigma_{x<t}$ AUROC
68.0
77.3
75.0
84.8
4.4 An online detection setting
Motivation. Just like image generative models typically have a filter to screen harmful generated contents [28], we believe building a similar mechanism for LLMs to detect memorized generated content in an online fashion would be helpful. Say, for example, that after generating a few sentences, the model for some reason suddenly starts to emit copyrighted or private information that is memorized from the training data. In such case, an effective mitigation would be to use pre-training data detection method internally to identify that a part of the generated text is training data, and then let the model refrain from further generation.
Setup. Current benchmarks fail to simulate the online detection setting, since each whole input is either pure training text or pure non-training text. To address this, we adapt WikiMIA to construct an online version of the benchmark. Concretely, each input text is created by concatenating a training text at the end of a non-training text, closely simulating the representative scenario discussed above. Both the training and non-training text have random length, varying among {32, 64, 128}. In this online setting, the prediction on each part of the input, instead of on the whole input, is of interests. Therefore, we split each input into chunks with a length of 32. Methods will be operating on all chunks, and the performance is evaluated on the chunk level. Essentially, this setup simulates using a non-overlapping sliding window of size 32 to sweep over the texts and detecting whether each part within the sliding window is training data or not. See Appendix C Figure 3 for visual illustration.
Table 4: AUROC results in the online detection setting with LLaMA models.
13B 30B 65B
Loss [39] 58.1 61.4 64.0
Zlib [4] 58.3 61.5 64.0
Min-K[29] 58.4 61.6 64.1
Min-K%++ 68.1 67.4 68.4
Results. Table 4 summarizes the results. Note that the Ref, Lowercase, and Neighbor method are no longer applicable or practical in online setting since they require repeated model inference. We see that Min-K%++ is still the most reliable method for online detection. The numbers are lower than those in the conventional offline setting (Table 1), which is expected because now we can only evaluate $\log p(x_{t}|x_{<t},y)$ with $y$ being the prepended text, rather than the exact $\log p(x_{t}|x_{<t})$.
5 Related Work
Membership Inference Attacks. MIA has long been a security and privacy-related topic. Given a target model and a target input, MIA aims to infer whether the input has been used to train the target model [30, 39]. It has been extensively studied in both vision [5, 6, 8, 40] and language domains [4, 36, 23]. Besides investigations on advanced methodologies, MIA also has profound implications in quantifying privacy risks [25], measuring memorization [7], helping with privacy auditing [31, 38], and detecting test-set contamination [27] and copyrighted contents [24, 11].
Pre-training data detection for LLMs. Despite being an instance of MIA (the general definition and evaluation metrics remain the same), this problem poses unique challenges compared with conventional MIA settings [29, 10]. For example, many early MIA methods need to train shadow models on the same distribution as the target model, which is no longer practical as many LLMs’ pre-training corpus is nonpublic. Second, characteristics of LLM pre-training (e.g., few training epochs, large-scale training) inherently makes MIA much more challenging. Lastly, most existing works on MIA against LLMs target the fine-tuning stage [26, 12, 23], which cannot transfer to the pre-training stage for the same reasons.
As a result, there has been growing interests in pre-training data detection recently, though it is still largely underexplored. Shi et al. [29], to our knowledge, is the first to investigate this problem. They contribute the WikiMIA benchmark and propose the Min-Kmethod. Duan et al. [10] perform systematic evaluation with the constructed MIMIR benchmark and analyze the challenges. In this work we propose a novel method that outperforms Min-Kand achieve superior performances over existing reference-free methods on both benchmarks. There are also works that specifically investigate copy-righted content detection, e.g., identifying memorized books [11].
6 Conclusion and Discussion
In this work, we propose Min-K%++ as a novel method for pre-training data detection for LLMs. Motivated by our insight that training data tends to be local maximum or locates near local maximum along input dimensions, we design our method to examine whether the input forms the mode or has relatively high probability under the conditional categorical distribution of the LLM. It consistently achieves superior performances on two existing benchmarks and in various settings, which establishes a solid baseline for future studies. We hope that our method, along with our theoretical and empirical analysis, can motivate more research upon the pre-training data detection problem.
Limitation. A minor limitation of Min-K%++ is that it cannot be applied to a pure black-box model or API where logits are not available (e.g., OpenAI API only returns at most top-4 token probabilities). However, we emphasize that this not a technical limitation since our method follows the standard grey-box access assumption [23, 29, 10] and is applicable to all open-source LLMs.
Future work. Even with the proposed Min-K%++, detection on MIMIR is still challenging given the limited distribution shifts between training and non-training data. Understanding why and how exactly distribution shifts affect detection remains an important direction for future study. Meanwhile, as aforementioned, our theoretical insight discussed in Section 3.1 has potential to generalize to other modalities and models. Investigating whether the concept of our method can work for other applications would also be meaningful.
Appendix
Appendix A Existing Assets
The two existing assets used in this work are the WikiMIA benchmark [29] and the MIMIR benchmark [10]. Please see their information below.
Appendix B Broader Impacts
As our work proposes an advanced pre-training data detection method, we will discuss societal impacts around general pre-training data detection techniques.
Positive impacts. Pre-training data detection has positive implications for several serious issues. For example, it can help with privacy auditing [29] by detecting whether a data instance has indeed been forgot or unlearned by the model. It can also help content creators or media see whether their copyrighted contents have been illegally used to train LLMs [14, 20].
Negative impacts. Pre-training data detection, as an instance of membership inference attack, may be thought to exhibit negative impacts such as stealing intellectual property (proprietary training data) and exposing the sensitive information encoded in training samples [39]. However, such risks are much less profound and practical than data extraction attacks [8]. For example, the proprietary training corpus of big companies will most likely be kept in-house, which is nearly impossible for attackers to access in the first place, not to mention to perform the subsequent membership inference attack. Therefore, from a practical perspective, we posit that the risks of pre-training data detection as an “attack” would not pose severe negative societal impacts.
Appendix C Additional Results
We present additional results here. Tables 5 and 6 show the AUROC results on WikiMIA with all the 10 models in the original and paraphrased setting, respectively. Tables 7 and 8 show the TPR at low FPR results. Table 9 shows the TPR at low FPR results on MIMIR. Lastly, Figure 3 shows a demo of the online detection setting that we have discussed in Section 4.4.
Table 5: Full AUROC results on WikiMIA in the original (verbatim training text) setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 61.0 62.2 60.9 61.9 64.1 63.2 66.8
Mamba-2.8B 64.1 67.0 63.6 64.7 67.0 66.1 69.3
Pythia-2.8B 61.4 61.3 60.9 62.1 64.2 61.8 64.4
Pythia-6.9B 63.8 63.6 62.2 64.3 65.8 66.3 70.3
Pythia-12B 65.4 65.1 64.8 65.8 66.6 68.1 72.3
NeoX-20B 68.8 67.2 68.0 69.0 70.2 71.8 75.0
LLaMA-13B 67.5 57.9 64.0 67.8 65.8 68.0 84.8
LLaMA-30B 69.4 63.5 64.1 69.8 67.6 70.1 84.3
LLaMA-65B 70.7 68.8 66.5 71.1 69.6 71.3 85.1
OPT-66B 65.7 68.7 63.0 66.0 68.2 67.7 70.2
[…]
Average 65.8 64.5 63.8 66.3 66.9 67.4 74.3
64 Mamba-1.4B 58.2 60.6 57.0 60.4 60.6 62.2 67.2
Mamba-2.8B 61.2 64.3 61.7 63.0 63.6 65.4 70.6
Pythia-2.8B 58.4 59.6 57.8 60.6 61.3 61.2 65.0
Pythia-6.9B 60.7 62.4 58.2 62.6 63.2 65.0 71.6
Pythia-12B 61.9 63.0 59.6 63.5 62.6 67.8 72.6
NeoX-20B 66.2 65.7 65.8 67.6 67.1 72.2 76.0
LLaMA-13B 63.6 63.4 62.0 65.3 64.1 66.0 85.7
LLaMA-30B 66.2 69.0 62.1 67.5 67.1 68.5 84.7
LLaMA-65B 67.9 73.4 64.5 69.1 69.6 69.8 83.8
OPT-66B 62.3 67.0 61.2 63.9 64.1 67.0 70.0
[…]
Average 62.7 64.8 61.0 64.4 64.3 66.5 74.7
128 Mamba-1.4B 63.3 62.0 58.5 65.6 64.8 66.8 68.8
Mamba-2.8B 66.2 66.9 62.4 68.5 67.7 71.0 73.4
Pythia-2.8B 62.8 59.6 59.5 65.0 65.2 66.8 66.8
Pythia-6.9B 65.1 63.3 60.5 67.6 67.5 69.5 70.7
Pythia-12B 65.8 63.9 61.4 67.8 67.1 70.7 72.7
NeoX-20B 70.1 67.8 67.7 71.8 71.6 75.0 75.9
LLaMA-13B 67.8 62.6 60.6 69.7 68.3 71.5 83.9
LLaMA-30B 70.3 71.9 59.1 71.8 72.2 73.9 82.6
LLaMA-65B 70.7 73.7 63.3 72.1 73.7 73.6 80.0
OPT-66B 65.5 66.9 59.3 67.5 67.7 70.5 72.3
[…]
Average 66.8 65.8 61.2 68.7 68.6 70.9 74.7
Table 6: Full AUROC results on WikiMIA in the paraphrased setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 61.4 62.3 60.6 62.3 63.6 62.9 66.1
Mamba-2.8B 64.5 66.6 63.5 64.8 66.3 65.3 67.9
Pythia-2.8B 61.6 61.2 60.3 62.3 64.5 61.7 62.4
Pythia-6.9B 64.1 63.5 61.7 64.2 65.5 65.2 68.0
Pythia-12B 65.6 64.9 64.4 65.9 66.8 67.2 69.8
NeoX-20B 68.2 66.3 66.7 68.2 68.3 69.7 69.6
LLaMA-13B 68.0 56.2 63.2 68.3 65.0 68.4 82.7
LLaMA-30B 70.2 62.4 61.2 70.4 66.3 70.7 81.2
LLaMA-65B 71.8 68.2 64.8 72.0 68.7 72.2 81.4
OPT-66B 65.3 68.2 62.7 65.4 66.7 66.3 68.1
[…]
Average 66.1 64.0 62.9 66.4 66.2 67.0 71.7
64 Mamba-1.4B 56.4 59.6 57.0 59.1 60.6 58.0 63.3
Mamba-2.8B 59.8 64.5 62.0 61.9 63.7 62.4 65.8
Pythia-2.8B 56.5 59.2 56.1 59.0 59.6 56.8 58.5
Pythia-6.9B 59.3 62.9 57.7 61.6 63.1 61.1 64.8
Pythia-12B 60.0 63.2 59.1 62.1 62.8 62.5 65.8
NeoX-20B 64.4 65.9 65.1 66.4 67.4 66.1 67.5
LLaMA-13B 63.1 60.9 61.0 65.3 64.7 64.0 78.8
LLaMA-30B 65.5 65.4 59.8 67.4 66.7 65.7 74.9
LLaMA-65B 67.7 71.0 61.9 69.3 69.5 67.9 74.0
OPT-66B 60.4 67.9 60.1 62.3 64.6 62.6 64.7
[…]
Average 61.3 64.0 60.0 63.4 64.3 62.7 67.8
128 Mamba-1.4B 62.7 61.1 57.7 65.3 62.6 64.5 65.6
Mamba-2.8B 65.7 66.6 61.2 68.3 64.6 68.0 70.0
Pythia-2.8B 62.3 59.5 59.6 65.0 61.9 64.7 63.4
Pythia-6.9B 64.7 62.9 60.0 67.4 64.3 67.0 66.8
Pythia-12B 65.4 63.9 60.4 67.9 64.3 68.5 68.8
NeoX-20B 69.5 67.8 67.4 71.8 69.6 72.6 72.2
LLaMA-13B 67.2 59.7 56.4 69.6 64.0 68.7 76.2
LLaMA-30B 69.2 70.0 55.4 71.5 67.2 70.2 73.8
LLaMA-65B 70.2 72.0 60.1 72.1 70.3 70.8 70.7
OPT-66B 64.5 66.8 57.4 66.9 63.4 67.2 68.2
[…]
Average 66.1 65.0 59.5 68.6 65.2 68.2 69.6
Table 7: Full TPR at low FPR (FPR=5%) results on WikiMIA in the original (verbatim training text) setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 14.2 7.8 11.1 15.5 11.9 14.7 12.9
Mamba-2.8B 14.7 9.8 16.8 16.3 16.0 18.1 13.4
Pythia-2.8B 14.7 6.2 11.1 15.8 15.0 17.1 14.2
Pythia-6.9B 14.2 6.7 10.6 16.3 16.5 17.8 17.1
Pythia-12B 17.1 9.0 16.3 17.1 19.4 23.0 18.6
NeoX-20B 19.9 15.5 18.1 19.9 22.2 27.9 19.4
LLaMA-13B 13.9 4.7 9.6 11.6 11.6 18.9 38.5
LLaMA-30B 18.4 9.8 11.4 14.5 9.3 21.2 31.3
LLaMA-65B 22.5 12.4 12.1 18.6 6.5 26.1 41.1
OPT-66B 14.2 10.8 10.6 16.0 21.7 22.0 19.4
[…]
Average 16.4 9.3 12.8 16.1 15.0 20.7 22.6
64 Mamba-1.4B 9.5 4.6 8.8 14.1 8.8 19.4 16.6
Mamba-2.8B 10.2 9.2 16.6 14.8 10.6 19.0 21.5
Pythia-2.8B 10.2 10.6 10.2 14.4 10.2 18.3 16.2
Pythia-6.9B 13.4 12.0 11.6 16.2 10.9 19.0 26.1
Pythia-12B 9.2 13.0 12.3 11.3 11.3 21.5 20.1
NeoX-20B 13.0 15.5 15.5 16.6 13.0 20.4 20.4
LLaMA-13B 11.3 4.2 11.6 12.7 10.2 17.2 34.1
LLaMA-30B 13.7 11.3 11.3 15.5 9.9 17.6 36.3
LLaMA-65B 15.1 13.0 12.3 16.9 9.9 18.0 38.4
OPT-66B 13.4 13.0 10.9 13.4 12.0 26.4 22.5
[…]
Average 11.9 10.6 12.1 14.6 10.7 19.7 25.2
128 Mamba-1.4B 11.5 10.1 12.9 19.4 15.8 16.6 16.6
Mamba-2.8B 19.4 10.1 13.7 23.7 15.1 25.9 21.6
Pythia-2.8B 9.3 10.1 10.8 18.7 8.6 13.7 17.3
Pythia-6.9B 14.4 13.7 12.9 20.9 10.8 18.0 22.3
Pythia-12B 18.0 12.2 12.9 23.7 10.1 25.2 20.9
NeoX-20B 18.7 15.8 12.2 23.0 15.8 25.2 23.0
LLaMA-13B 21.6 10.8 15.8 18.7 12.9 25.9 43.2
LLaMA-30B 23.7 10.8 10.1 18.0 15.1 23.7 40.3
LLaMA-65B 23.0 18.0 14.4 22.3 15.8 23.7 27.3
OPT-66B 20.9 17.3 14.4 21.6 12.9 23.0 16.6
[…]
Average 18.1 12.9 13.0 21.0 13.3 22.1 24.9
Table 8: Full TPR at low FPR (FPR=5%) results on WikiMIA in the paraphrased setting.
Len. Model Loss Ref Lowercase Zlib Neighbor Min-KMin-K%++
32 Mamba-1.4B 14.2 5.9 13.2 13.2 7.2 15.2 10.6
Mamba-2.8B 16.5 10.1 15.0 12.7 9.3 19.9 13.4
Pythia-2.8B 14.2 7.2 11.6 14.5 8.5 16.5 13.9
Pythia-6.9B 15.0 6.2 11.9 12.7 9.6 21.7 17.1
Pythia-12B 17.3 8.0 16.5 15.5 9.8 19.9 17.3
NeoX-20B 18.1 15.2 15.5 18.6 15.2 19.6 12.9
LLaMA-13B 16.3 5.4 9.6 15.0 8.5 17.6 35.9
LLaMA-30B 14.7 7.5 12.7 15.0 9.3 18.1 27.4
LLaMA-65B 23.3 9.3 11.9 16.5 12.1 24.3 35.9
OPT-66B 15.2 10.3 13.4 17.1 12.1 18.1 15.2
[…]
Average 16.5 8.5 13.1 15.1 10.2 19.1 20.0
64 Mamba-1.4B 8.1 8.1 9.5 15.1 9.5 8.4 7.0
Mamba-2.8B 12.3 11.3 14.8 14.8 18.3 13.0 12.3
Pythia-2.8B 9.5 13.0 11.3 16.6 11.3 11.3 9.9
Pythia-6.9B 10.6 16.2 11.3 15.8 12.7 12.7 14.1
Pythia-12B 11.6 14.4 13.4 16.2 10.6 14.4 13.7
NeoX-20B 16.2 14.1 13.7 19.4 18.3 17.6 13.0
LLaMA-13B 12.0 4.6 13.7 13.4 14.4 13.4 26.4
LLaMA-30B 13.4 8.1 8.1 16.9 11.6 14.4 21.5
LLaMA-65B 13.4 10.9 9.5 18.0 16.9 13.7 29.2
OPT-66B 13.4 13.0 13.4 14.8 13.7 14.8 12.7
[…]
Average 12.0 11.4 11.9 16.1 13.7 13.4 16.0
128 Mamba-1.4B 13.7 11.5 11.5 17.3 13.7 14.4 10.1
Mamba-2.8B 16.6 10.8 15.1 20.1 17.3 20.1 15.1
Pythia-2.8B 14.4 7.2 8.6 16.6 12.2 14.4 14.4
Pythia-6.9B 16.6 8.6 11.5 20.9 17.3 17.3 21.6
Pythia-12B 19.4 8.6 12.2 19.4 10.1 21.6 17.3
NeoX-20B 15.8 19.4 16.6 21.6 18.7 22.3 19.4
LLaMA-13B 18.0 4.3 15.8 21.6 13.7 20.1 35.2
LLaMA-30B 18.7 18.7 13.7 19.4 14.4 18.7 21.6
LLaMA-65B 24.5 12.9 13.7 22.3 18.7 25.2 25.2
OPT-66B 18.0 15.8 11.5 18.7 12.9 20.1 18.7
[…]
Average 17.6 11.8 13.0 19.8 14.9 19.4 19.9
Table 9: TPR at low FPR (FPR=5%) results on MIMIR. In each column, the best result across all methods is bolded, with the runner-up underlined. Ref relies on a reference LLM. Neighbor induces significant extra computational cost than others (25×25$\times$25 × in this case), for which reason we don’t run on the 12B model.
Wikipedia Github Pile CC PubMed Central
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [39] 4.2 4.7 4.7 5.1 5.0 22.6 32.1 33.6 38.5 40.7 3.1 5.0 4.8 4.9 5.1 4.0 4.4 4.3 4.9 5.0
Ref [4] 6.1 5.3 5.5 5.6 5.7 23.4 14.8 14.9 15.4 16.2 5.5 5.6 5.8 5.8 7.5 5.7 4.1 4.0 5.9 8.7
Zlib [4] 4.2 5.7 5.9 6.3 6.8 25.0 32.8 36.1 39.3 40.8 4.0 5.1 5.4 6.2 6.6 3.8 3.6 3.5 4.3 4.4
Neighbor [23] 4.0 4.5 4.9 5.8 / 24.7 31.6 29.8 34.1 / 3.9 3.6 4.0 5.3 / 3.9 3.7 4.5 4.5 /
Min-K[29] 6.4 5.6 6.4 6.5 8.1 23.3 32.2 34.0 39.0 40.8 4.2 5.1 5.2 5.5 5.7 4.7 5.2 4.8 5.9 5.4
Min-K%++ 5.7 6.1 8.5 11.4 11.5 25.4 33.2 34.2 38.2 40.1 5.8 5.0 5.4 6.3 6.3 5.1 6.3 6.5 7.4 9.0
ArXiv DM Mathematics HackerNews Average
Method 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B 160M 1.4B 2.8B 6.9B 12B
Loss [39] 4.0 4.8 4.6 5.4 5.6 3.8 4.3 4.1 4.1 4.0 5.0 4.8 5.5 5.9 6.8 6.7 8.6 8.8 9.8 10.3
Ref [4] 5.2 5.4 5.9 6.8 7.4 5.3 3.6 4.9 5.4 5.9 5.2 6.3 7.6 7.0 7.6 8.1 6.4 6.9 7.4 8.4
Zlib [4] 2.9 4.3 4.1 4.6 4.7 4.1 5.0 4.6 4.3 4.3 5.0 5.5 5.8 5.6 5.8 7.0 8.9 9.3 10.1 10.5
Neighbor [23] 4.7 4.8 4.4 4.1 / 5.6 4.4 4.5 4.5 / 6.5 5.2 5.3 5.7 / 7.6 8.3 8.2 9.1 /
Min-K[29] 4.9 4.8 4.7 5.6 6.2 4.5 4.5 4.6 4.7 5.2 5.2 5.7 5.9 6.3 6.9 7.6 9.0 9.4 10.5 11.2
Min-K%++ 6.0 6.0 6.7 8.2 8.6 5.4 5.5 5.7 6.2 6.3 5.5 4.9 5.7 6.6 6.6 8.4 9.6 10.4 12.0 12.6
Figure 3: A demo of Min-K%++ in the online detection setting. Each input is split into 32-length chunks or sliding windows (separated with the mark ‘|’), and Min-K%++ performs detection over each chunk. The token-wise scores that contribute to the chunk-wise score are marked by the darker colors. The detection results are marked by non-training data v.s. training data. During LLM’s generation, once we detect training data in the current chunk we can stop further generation to avoid potential issues.